### v1

# thanks: ORCID: $0$-0007-0014-2446

Abstract
We study the evolution of a spectral deformation flow for amplitudes $C_{n}(\tau)$, governed by a second-order self-adjoint operator $\hat{C}$ on a compact domain. This flow leads to exponential stabilization $C_{n}(\tau)\to\pi$, encoded in the spectral multi-function $C(v,\tau,n)$, which describes the geometric deformation across scales. Building on prior results on spectral rigidity, completeness, and asymptotics of $\hat{C}$, we analyze conditions under which this convergence determines the full topological and smooth structure of simply-connected closed manifolds in all dimensions $d\geq 3$. Using spectral asymptotics and global analysis, we show that such convergence uniquely selects the standard sphere, excluding exotic smooth structures and isospectral non-isometric manifolds within the operator domain. In particular, we examine the critical case $d=4$ and prove that no non-standard smooth structure can yield the limiting spectral profile. These results suggest that the spectrum of $\hat{C}$, shaped by its geometric flow, serves as a complete invariant of smooth geometry.
I Introduction
The deformation function $C(v)=\pi(1-v^{2}/c^{2})$, introduced in  alexa2025-flow , is a real-analytic, even function originally defined on the open interval $v\in(-c,c)$, where $c$ denotes the speed of light, and the function satisfies $C(\pm c)=0$, $C(0)=\pi$. To restrict the deformation dynamics to geometrically admissible configurations and ensure analytic control, we introduced a cutoff scale — the critical velocity $v_{c}:=\sqrt{1-\tfrac{1}{\pi}}\cdot c\approx 0.8257\,c$ — defined by the threshold condition $C(v_{c})=1$. This selects the compact interval $v\in[-v_{c},v_{c}]$, on which the deformation profile satisfies $C(v)\geq 1$, remains strictly positive and bounded, and excludes configurations of extreme curvature. This domain enables the formulation of a globally regular scalar flow $C(v,\tau)$, governed by a scalar parabolic variational equation, that remains globally regular and avoids the curvature singularities characteristic of tensorial Ricci flows hamilton1982 ; perelman2002 , where singularity formation requires surgery and blow-up analysis. Within this framework, we constructed a deformation energy functional, established global convergence of the flow, and rigorously demonstrated that any simply-connected compact 3-manifold in the conformal-scalar class evolves asymptotically to the standard 3-sphere under the flow.
Then we introduced alexa2025-operator a second-order differential operator $\hat{C}:=\pi\left(1+\frac{\hbar^{2}}{c^{2}}\frac{d^{2}}{dv^{2}}\right)$, acting on the Hilbert space $L^{2}([-v_{c},v_{c}])$ with Dirichlet boundary conditions and Sobolev domain $H^{2}\cap H^{1}_{0}$. We proved that $\hat{C}$ is essentially self-adjoint by direct computation of von Neumann’s deficiency indices, which vanish identically. This established the mathematical consistency of the operator and ensured a well-defined spectral problem on compact domains. The resulting operator encodes quantized geometric modes of deformation and provides a rigorous foundation for subsequent spectral analysis.
Building on this, the third article alexa2025-spectrum developed a full spectral theory of the operator $\hat{C}$, including the proof of completeness of its eigenfunctions, a sharp asymptotic estimate $C_{n}\sim\pi-\kappa n^{2}$, and the derivation of the Spectral Rigidity Theorem. This theorem asserts that the limiting constant spectrum $C_{n}=\pi$ corresponds uniquely (up to isometry) to the standard sphere. We also established that the eigenbasis $\{\psi_{n}\}$ forms a complete orthonormal system in $L^{2}([-v_{c},v_{c}])$, and that the geometric deformation $C(v)$ admits a unique spectral reconstruction in terms of its expansion over the modes $\psi_{n}$. In this way, the spectrum of $\hat{C}$ encodes the full geometric content of the deformation profile.
In the present work, we investigate the asymptotic behavior of the spectral deformation flow $C_{n}(\tau)$, where the amplitudes evolve under a variational evolution derived from the geometry. The full geometric content of this flow is encoded in the spectral multi-function $C(v,\tau,n):=C_{n}(\tau)\psi_{n}(v)$, which resolves the deformation profile simultaneously across scale, spectral mode, and time. Our goal is to characterize the class of manifolds for which the flow stabilizes to the symmetric configuration $C_{n}(\tau)\to\pi$, and to determine whether such convergence uniquely determines both the topological and smooth structure of the underlying space. Building on the spectral theory of the operator $\hat{C}$, we prove that exponential convergence of the amplitudes implies smooth geometric convergence in the $C^{\infty}$ topology, enforced by the saturation of all spectral invariants—including the Weyl coefficients, the $\eta$-invariant, and the analytic torsion. We establish global stability of the flow, derive entropy bounds and scaling laws, and show that no exotic smooth structures or isospectral non-isometric counterparts can yield the limiting spectrum within the operator domain. In particular, we examine the critical four-dimensional case $d=4$, and demonstrate that the spectral attractor $C_{n}(\tau)\to\pi$ cannot be produced by any non-standard smooth structure on $S^{4}$. These results extend the spectral classification program initiated in alexa2025-flow ; alexa2025-operator ; alexa2025-spectrum , and demonstrate that the geometric flow of spectral amplitudes—encoded in the multi-function $C(v,\tau,n)$—acts as a complete invariant of smooth geometry in all dimensions $d\geq 3$.
II Spectral Geometry and the Multifunction $C(v,\tau,n)$
In this section, we construct the spectral framework that underlies the deformation geometry defined by the function $C(v,\tau,n)$. Building upon the self-adjoint operator introduced in alexa2025-operator , and the spectral decomposition analyzed in alexa2025-spectrum , we define a canonical multifunction on a Hilbert space of square-integrable functions over a compact interval. This multifunction serves as a universal representation of geometric deformation modes and forms the basis for the spectral flow and topological classification developed in later sections.
We begin by recalling the operator $\hat{C}$ as a second-order differential operator encoding the relativistic deformation of geometry. We then establish its discrete spectrum and the associated orthonormal eigenbasis, providing the functional foundation for defining the geometric state space. This basis enables the expansion of deformations into elementary spectral modes, and allows us to formulate a dynamic evolution directly in the spectral domain.
II.1 Spectral Operator and Eigenbasis
Let $C(v)=\pi\left(1-\frac{v^{2}}{c^{2}}\right)$ denote the relativistic deformation profile defined over the compact interval $v\in[-v_{c},v_{c}]$, with the critical velocity bound $v_{c}=c\sqrt{1-1/\pi}$ introduced in alexa2025-flow . To quantize this scalar geometry, we introduce the second-order differential operator
$\hat{C}:=\pi\left(1+\frac{\hbar^{2}}{c^{2}}\frac{d^{2}}{dv^{2}}\right),$ (1)
acting on the domain $D(\hat{C}):=H^{2}([-v_{c},v_{c}])\cap H^{1}_{0}([-v_{c},v_{c}])$, that is, the Sobolev space of twice weakly differentiable functions vanishing at the endpoints. As shown in alexa2025-operator , the operator $\hat{C}$ is essentially self-adjoint on this domain, with deficiency indices (0,0)(0,0)( 0 , 0 ), satisfying the von Neumann criterion for self-adjointness of symmetric operators on a compact interval.
The spectral theory of Sturm–Liouville operators on compact domains (see zettl2005 , reed-simon4 ) ensures that $\hat{C}$ admits a purely discrete spectrum with real, simple eigenvalues and a complete orthonormal set of eigenfunctions in $L^{2}([-v_{c},v_{c}])$. The eigenvalue problem
$\hat{C}\psi_{n}(v)=C_{n}\psi_{n}(v),\qquad\psi_{n}(\pm v_{c})=0,$ (2)
has solutions indexed by $n\in\mathbb{N}_{0}$, forming a countable orthonormal basis $\{\psi_{n}\}_{n=0}^{\infty}$. These eigenfunctions satisfy
$\langle\psi_{n},\psi_{m}\rangle=\delta_{nm},\qquad\sum_{n=0}^{\infty}\psi_{n}(v)\psi_{n}(v^{\prime})=\delta(v-v^{\prime}).$ (3)
This holds in the distributional sense.
The corresponding eigenvalues $C_{n}\in\mathbb{R}$ are strictly decreasing, with asymptotic behavior derived in alexa2025-spectrum :
$C_{n}=\pi-\frac{\pi^{3}\hbar^{2}}{4c^{2}v_{c}^{2}}(n+1)^{2}+o(n^{2}),\qquad n\to\infty.$ (4)
This spectrum is strictly bounded above by $\pi$ and unbounded below, reflecting the ultraviolet spectral rigidity of the operator. The simplicity of the spectrum follows from the classical oscillation theorems for Dirichlet Sturm–Liouville problems zettl2005 , and guarantees that the eigenfunctions $\psi_{n}$ are uniquely defined up to sign.
For any admissible deformation profile $C(v)\in D(\hat{C})$, the spectral decomposition
$C(v)=\sum_{n=0}^{\infty}C_{n}\psi_{n}(v),$ (5)
converges in $L^{2}([-v_{c},v_{c}])$, and, for smooth profiles, in the $C^{\infty}$-topology as well, due to the ellipticity of the operator and standard spectral regularity theorems (see reed-simon4 , Vol. IV).
Moreover, for every integer $k\geq 0$, the eigenfunctions $\psi_{n}$ satisfy the uniform derivative bounds
Lemma 1 (Derivative Bounds).
For each $k\in\mathbb{N}$, there exists a constant $C_{k}>0$ such that
$\sup_{v\in[-v_{c},v_{c}]}\left|\frac{d^{k}}{dv^{k}}\psi_{n}(v)\right|\leq C_{k}(n+1)^{k+1},\qquad\forall n\in\mathbb{N}_{0}.$ (6)
Proof.
This estimate follows from the explicit sine form of the eigenfunctions:
$\psi_{n}(v)=\sqrt{\frac{1}{v_{c}}}\sin\left(\frac{(n+1)\pi}{2v_{c}}(v+v_{c})\right),$ (7)
whose $k$-th derivative scales as $\sim(n+1)^{k}$, and the uniform $L^{\infty}$-bound on the interval yields the extra factor, giving $(n+1)^{k+1}$. A detailed derivation is provided in alexa2025-spectrum , Lemma 4. ∎
These estimates, derived in alexa2025-spectrum , are essential for controlling nonlinear spectral interactions in the variational flow constructed in Section III.
This spectral framework provides a canonical Hilbert basis for encoding geometric deformation. In the next subsection, we use this structure to define the multifunction $C(v,\tau,n)$, representing evolving scalar geometry in terms of spectral amplitudes and eigenmodes. This decomposition forms the foundation for the variational dynamics and topological classification that follow.
II.2 Definition of the Multifunction
Having established the existence of a discrete orthonormal basis $\{\psi_{n}(v)\}$ of eigenfunctions of the operator $\hat{C}$, we now define the central object of this work — a spectral multifunction that unifies geometric structure and spectral dynamics. For each fixed mode index $n\in\mathbb{N}$, we define the deformation component by $C(v,\tau,n):=C_{n}(\tau)\,\psi_{n}(v)$, where the spectral amplitude $C_{n}(\tau)$ evolves under a flow parameter $\tau\geq 0$, while the eigenfunction $\psi_{n}(v)$ remains fixed. This construction reflects a separation between the static spectral geometry encoded in the operator $\hat{C}$, and the dynamical evolution of the geometry as encoded in the amplitudes $C_{n}(\tau)$.
Definition 1 (Spectral Multifunction).
The spectral multifunction $C(v,\tau,n):=C_{n}(\tau)\psi_{n}(v)$ is a family of time-dependent geometric components indexed by $n\in\mathbb{N}_{0}$, where $\{\psi_{n}\}$ is the eigenbasis of the operator $\hat{C}$, and $\{C_{n}(\tau)\}\in\ell^{2}(\mathbb{N}_{0})$ evolve under a variational flow.
The quantity $C(v,\tau,n)$ is not a single-valued function in the traditional sense, but a multifunction in the spectral domain — that is, an indexed family of time-evolving geometric components, each associated with a specific eigenmode of the operator. The total geometric profile at time $\tau$ is recovered via spectral synthesis as
$C(v,\tau)=\sum_{n=0}^{\infty}C_{n}(\tau)\psi_{n}(v),$ (8)
with convergence in the $L^{2}$-norm on the interval $[-v_{c},v_{c}]$, and under regularity assumptions, in the $C^{\infty}$-topology by standard results on spectral expansions reed-simon4 ; zettl2005 . Thus, the evolution of geometry is no longer expressed through local differential changes of a metric tensor, but through the variation of spectral amplitudes over a fixed, self-adjoint spectral basis.
This shift in perspective—from coordinate-based metric deformation to infinite-dimensional spectral dynamics—is more than notational. It reflects a fundamental principle: the geometry is entirely encoded in the spectrum and its evolution. The multifunction $C(v,\tau,n)$ thus provides a canonical spectral quantization of scalar geometry. The coordinate $v$ serves as a spectral parameter related to relativistic deformation, the index $n$ labels quantized geometric modes, and the flow parameter $\tau$ governs the dynamical evolution of the structure. In this setting, the geometric data becomes a trajectory in the Hilbert space $\ell^{2}(\mathbb{N}_{0})$, where each configuration $\{C_{n}(\tau)\}$ corresponds to a smooth deformation profile.
It is essential to note that this framework does not rely on any tensorial or manifold coordinates beyond the initial spectral interval. The spectrum of $\hat{C}$, together with its eigenfunctions, determines the admissible configurations entirely. The resulting evolution is governed by a variational principle which we formulate in the next section, leading to a spectral flow that drives all modes toward the symmetric configuration $C_{n}(\tau)\to\pi$, corresponding to the maximally symmetric geometry.
II.3 Spectral Space and Geometric Interpretation
The spectral decomposition described above defines a canonical representation of smooth deformation profiles as elements of the infinite-dimensional Hilbert space $\ell^{2}(\mathbb{N}_{0})$. Each admissible configuration $C(v,\tau)\in L^{2}([-v_{c},v_{c}])$, when expressed through the eigenbasis $\{\psi_{n}(v)\}$, corresponds to a unique sequence of spectral amplitudes $\{C_{n}(\tau)\}$, satisfying $\sum_{n}|C_{n}(\tau)|^{2}<\infty$. The collection of all such sequences forms the configuration space for the spectral dynamics of geometry. In this framework, geometric evolution is recast as a flow in $\ell^{2}$, and all physical observables are interpreted as functionals on this space.
The interpretation of $\ell^{2}$ as a space of geometric states is rooted in the completeness of the spectral basis and the self-adjointness of the deformation operator $\hat{C}$. Since the eigenfunctions $\psi_{n}(v)$ span $L^{2}([-v_{c},v_{c}])$, the full content of the geometry is encoded in the amplitudes $C_{n}$. In particular, the spectral energy, entropy, and higher invariants, as introduced in later sections, are naturally expressed as quadratic or logarithmic functionals on $\ell^{2}$. This parallels the formal structure of quantum field theory, where the configuration space is a function space over spectral modes, but in our case the modes emerge from geometric deformation, not from harmonic oscillators or particles.
Moreover, the convergence $C_{n}(\tau)\to\pi$ in $\ell^{2}$ norm plays a fundamental role in topological classification. It implies uniform geometric stabilization, and—under spectral rigidity—determines the diffeomorphism type of the underlying manifold. In this sense, the Hilbert space $\ell^{2}$ not only encodes the geometry but also becomes the stage on which topology is dynamically selected. This is in contrast to classical approaches, where topology is a fixed input. Here, topology is an emergent consequence of the spectral dynamics.
The geometric interpretation of the multifunction $C(v,\tau,n)$ is therefore nonlocal and intrinsically spectral. The variable $v$ is not a spatial coordinate on a manifold, but a spectral deformation parameter associated with a relativistic scale; it serves as the argument of the eigenfunctions $\psi_{n}(v)$ of the operator $\hat{C}$. The index $n$ labels intrinsic geometric excitations, and the flow parameter $\tau$ governs the dynamical evolution of the spectral amplitudes. What emerges from this structure is not a classical metric tensor, but a dynamically evolving spectral profile whose limiting behavior encodes the smooth and topological classification of the underlying space.
III Variational Spectral Flow and Global Stability
The spectral representation introduced in Section II recasts the geometry of deformation into a sequence of spectral amplitudes $\{C_{n}(\tau)\}_{n\in\mathbb{N}_{0}}$, evolving over a fixed self-adjoint eigenbasis $\{\psi_{n}(v)\}$ in the Hilbert space $L^{2}([-v_{c},v_{c}])$. This encoding allows for a natural formulation of geometry as a trajectory in $\ell^{2}$, where each configuration represents a smooth deformation profile. In this section, we define a variational principle that governs the evolution of these amplitudes and derive the corresponding flow equations. The formulation is motivated by the geometric idea that the most symmetric configuration $C_{n}=\pi\ \forall n$ minimizes a global energy functional, and that deviations from this state represent spectral asymmetry or geometric anisotropy.
The flow is constructed to evolve each amplitude $C_{n}(\tau)$ toward its maximal symmetric value $\pi$, while allowing for nonlinear interactions that preserve the overall variational structure. We derive a well-defined dynamical system in $\ell^{2}$, prove global well-posedness and exponential convergence, and establish that the variational evolution asymptotically stabilizes the geometry under spectral rigidity. This flow plays a role analogous to Ricci flow in Riemannian geometry, but operates directly in the spectral domain, with no reliance on local curvature tensors or manifold coordinates hamilton1982 .
III.1 Action Principle and Flow Equation
Let $\{C_{n}(\tau)\}_{n\in\mathbb{N}_{0}}$ denote the spectral amplitudes associated with the evolving deformation profile $C(v,\tau)\in L^{2}([-v_{c},v_{c}])$, with orthonormal basis $\{\psi_{n}(v)\}$ and fixed operator $\hat{C}$ as defined in Section II and detailed in alexa2025-operator . We introduce a variational structure by postulating that the evolution of $C_{n}(\tau)$ arises from the minimization of a global action functional defined on trajectories in $\ell^{2}$. The variational principle is defined via a Lagrangian functional ($\mathcal{L}(\{C_{n}(\tau)\},\{\dot{C}_{n}(\tau)\})$, and the corresponding Euler–Lagrange equations define the flow reed-simon1 .
We define the Lagrangian at each instant $\tau$ by:
$\mathcal{L}(\tau):=\sum_{n=0}^{\infty}\left[\frac{1}{2}\dot{C}_{n}^{2}(\tau)-\frac{\alpha_{n}}{2}(C_{n}(\tau)-\pi)^{2}\right],$ (9)
where $\dot{C}_{n}:=\frac{dC_{n}}{d\tau}$, and $\alpha_{n}>0$ is a fixed spectral weight that controls the stiffness of each mode. The term $(C_{n}-\pi)^{2}$ penalizes deviations from the maximally symmetric configuration, while the kinetic term $\dot{C}_{n}^{2}$ captures the spectral dynamics. This Lagrangian is quadratic and defines a convex action on the Hilbert space $\ell^{2}$, provided $\sum_{n}\alpha_{n}<\infty$, ensuring that high-frequency modes are properly controlled ambrosio2008 .
The corresponding action functional is given by:
[…]
and the variational principle requires that this action be stationary under variations $\delta C_{n}(\tau)$ with compact support in $\tau$. The Euler–Lagrange equation for each mode is then:
$\frac{d^{2}C_{n}}{d\tau^{2}}+\alpha_{n}(C_{n}-\pi)=0.$ (11)
Equation (11) describes a collection of decoupled harmonic oscillators with equilibrium position at $\pi$. The general solution of each component is:
$C_{n}(\tau)=\pi+A_{n}\cos(\sqrt{\alpha_{n}}\tau)+B_{n}\sin(\sqrt{\alpha_{n}}\tau),$ (12)
with constants $A_{n},B_{n}\in\mathbb{R}$ determined by initial conditions. However, in order to obtain monotonic convergence to $C_{n}=\pi$, we impose a gradient flow structure by introducing a frictional term, thereby replacing the conservative dynamics (11) with a dissipative flow evans-pde .
We define the first-order gradient flow associated with the functional:
$\mathcal{E}[C(\tau)]:=\sum_{n=0}^{\infty}\frac{\alpha_{n}}{2}(C_{n}(\tau)-\pi)^{2},$ (13)
and prescribe the evolution equations as:
$\frac{dC_{n}}{d\tau}=-\alpha_{n}(C_{n}-\pi).$ (14)
This system defines an autonomous linear ODE in $\ell^{2}$, and the solution is:
$C_{n}(\tau)=\pi+(C_{n}(0)-\pi)e^{-\alpha_{n}\tau}.$ (15)
It follows that $C_{n}(\tau)\to\pi$ exponentially as $\tau\to\infty$, with rate determined by $\alpha_{n}$. The system is globally well-posed and converges to the symmetric fixed point in $\ell^{2}$, provided that the initial sequence $\{C_{n}(0)\}\in\ell^{2}$ reed-simon1 . This linear gradient flow provides the foundation for analyzing more general nonlinear deformations introduced in the next subsection.
We note that the energy functional (13) is strictly decreasing along solutions of (14), as:
$\frac{d\mathcal{E}}{d\tau}=\sum_{n=0}^{\infty}\alpha_{n}(C_{n}-\pi)\frac{dC_{n}}{d\tau}=-\sum_{n=0}^{\infty}\alpha_{n}^{2}(C_{n}-\pi)^{2}\leq 0.$ (16)
This ensures that the flow is energetically dissipative and monotonic, and that no oscillatory behavior occurs. In particular, the fixed point $C_{n}=\pi$ is globally attractive in $\ell^{2}$, and represents a state of maximal spectral symmetry.
This variational formulation provides a rigorous basis for interpreting geometric evolution as a spectral relaxation toward symmetry. In the next subsection, we extend this linear system by introducing entropy, cubic interactions, and nonlinear coupling, to reflect more realistic geometric dynamics while preserving global stability.
III.2 Energy Functional and Entropy Decay
To formalize the dissipative structure of the spectral flow and quantify its irreversibility, we introduce two complementary functionals: the spectral energy and the geometric entropy. These functionals serve to characterize the deviation from maximal symmetry and encode the information-theoretic and variational content of the evolving geometry.
The spectral energy functional, already introduced in Eq. (13), is given by
$\mathcal{E}[\tau]:=\sum_{n=0}^{\infty}\frac{\alpha_{n}}{2}\left(C_{n}(\tau)-\pi\right)^{2},$ (17)
where $\alpha_{n}>0$ is a spectral stiffness coefficient that controls the rate of convergence of each mode. This functional attains its minimum value $\mathcal{E}[\tau]=0$ if and only if $C_{n}(\tau)=\pi\ \forall n$, which corresponds to the maximally symmetric geometric configuration. The convexity and coercivity of $\mathcal{E}$ on $\ell^{2}$ follow from the positivity of $\alpha_{n}$, and the functional defines a quadratic energy landscape in the space of spectral amplitudes.
The rate of decay of this energy along the flow is given by the time derivative:
$\frac{d\mathcal{E}}{d\tau}=-\sum_{n=0}^{\infty}\alpha_{n}^{2}\left(C_{n}(\tau)-\pi\right)^{2}=-2\sum_{n=0}^{\infty}\alpha_{n}\left(\frac{dC_{n}}{d\tau}\right)^{2}\leq 0,$ (18)
which confirms that the flow is strictly dissipative in the energy functional and monotonic with respect to spectral symmetry. No oscillatory or limit cycle behavior is possible in this linear regime. This structure generalizes the Lyapunov function formulation in dynamical systems theory to an infinite-dimensional setting hale1980 .
To capture the statistical structure of the spectral distribution and its convergence to order, we define the spectral entropy functional:
$\mathcal{S}[\tau]:=-\sum_{n=0}^{\infty}w_{n}(\tau)\log w_{n}(\tau),\qquad w_{n}(\tau):=\frac{|C_{n}(\tau)|^{2}}{\sum_{k}|C_{k}(\tau)|^{2}},$ (19)
which interprets the normalized squared amplitudes $w_{n}(\tau)\in[0,1]$ as a probability distribution over modes. This definition is standard in spectral signal theory and quantum information nielsen-chuang , and characterizes the disorder or uncertainty of the geometric state encoded in the spectrum.
As the flow evolves, each amplitude $C_{n}(\tau)$ tends toward the symmetric value $\pi$, and the normalized spectral distribution
$w_{n}(\tau)=\frac{|C_{n}(\tau)|^{2}}{\sum_{k}|C_{k}(\tau)|^{2}}$ (20)
becomes increasingly concentrated near equilibrium. Let $N(\tau,\varepsilon):=\#\{n:|C_{n}(\tau)-\pi|>\varepsilon\}$ denote the number of modes significantly deviating from the symmetric configuration. Then, as $\tau\to\infty$, we have $N(\tau,\varepsilon)\to 0$, and the spectral entropy
$\mathcal{S}[\tau]:=-\sum_{n}w_{n}(\tau)\log w_{n}(\tau)$ (21)
satisfies $\mathcal{S}[\tau]\sim\log N(\tau,\varepsilon)$, quantifying the effective complexity of the deformation. In the asymptotic regime, $\mathcal{S}[\tau]$ approaches a finite limit: it vanishes if all modes converge exactly to $\pi$, or stabilizes at a positive constant determined by the number of persistent modes with $|C_{n}(\tau)-\pi|\neq 0$. This decay reflects the geometric ordering induced by the flow and parallels thermodynamic entropy reduction in dissipative systems.
In particular, for initial data $\{C_{n}(0)\}\in\ell^{2}$ satisfying uniform spectral decay, the entropy functional $\mathcal{S}[\tau]$ is finite and differentiable for all $\tau>0$, and satisfies:
$\frac{d\mathcal{S}}{d\tau}=-\sum_{n=0}^{\infty}\left(\frac{2C_{n}\dot{C}_{n}}{\|C\|^{2}}-\frac{2C_{n}^{2}}{\|C\|^{4}}\sum_{k}C_{k}\dot{C}_{k}\right)\log w_{n},$ (22)
where $\|C\|^{2}:=\sum_{n}C_{n}^{2}$. This expression is negative definite in the asymptotic regime $C_{n}(\tau)\to\pi$, confirming that entropy strictly decreases along the flow.
The dual decay of spectral energy and entropy formalizes the irreversible geometric relaxation and defines a spectral arrow of time. Unlike in the Ricci flow, where entropy arises from curvature functionals (e.g., Perelmans entropy perelman2002 ), here entropy is directly tied to the statistical structure of spectral deformation. This provides a natural information-theoretic interpretation of geometric evolution in terms of spectral compression and stabilization.
III.3 Cubic Interactions and Nonlinear Coupling
While the linear gradient flow introduced in the previous subsection ensures exponential convergence to the symmetric configuration $C_{n}=\pi$, it neglects the possible interactions between spectral modes that arise from nonlinear geometry. To incorporate these effects, we extend the linear model by introducing cubic interactions that preserve the variational structure while coupling different modes through fixed interaction coefficients. These terms are analogous to interaction vertices in quantum field theory, but they arise here from the nonlinear structure of geometric deformation.
We define the nonlinear evolution equations by introducing a cubic coupling term:
$\frac{dC_{n}}{d\tau}=-\alpha_{n}(C_{n}-\pi)+\sum_{k,m}g_{nkm}C_{k}(\tau)C_{m}(\tau),$ (23)
where the coefficients $g_{nkm}\in\mathbb{R}$ encode the strength of interaction between modes $k$, $m$, and $n$. These coefficients are assumed to be symmetric in $k$ and $m$, and decay sufficiently fast in all indices to ensure convergence of the series in (23). A natural choice is
$|g_{nkm}|\leq\frac{C}{(1+n^{p}+k^{p}+m^{p})},$ (24)
for some constants $C>0$ and $p>2$, ensuring absolute convergence of the triple sum for sequences in $\ell^{2}$, and preserving the Hilbert space structure.
The flow (23) defines an infinite-dimensional dynamical system with nonlinear coupling, and can be interpreted as the gradient flow of a perturbed energy functional
$\mathcal{E}[C]=\sum_{n=0}^{\infty}\frac{\alpha_{n}}{2}(C_{n}-\pi)^{2}-\sum_{n,k,m}\frac{1}{3}g_{nkm}C_{n}C_{k}C_{m},$ (25)
which is a cubic perturbation of the quadratic energy introduced earlier. The evolution equation (23) satisfies
$\frac{dC_{n}}{d\tau}=-\frac{\delta\mathcal{E}}{\delta C_{n}},$ (26)
where the functional derivative is understood in the weak $\ell^{2}$-topology. This guarantees that the flow decreases the perturbed energy:
$\frac{d\mathcal{E}}{d\tau}=\sum_{n}\frac{\delta\mathcal{E}}{\delta C_{n}}\frac{dC_{n}}{d\tau}=-\sum_{n}\left(\frac{dC_{n}}{d\tau}\right)^{2}\leq 0,$ (27)
so long as the flow remains well-posed in $\ell^{2}$.
Existence and uniqueness of solutions for (23) with initial data $\{C_{n}(0)\}\in\ell^{2}$ follow from standard Picard iteration theorems in Banach spaces, using that the nonlinear term is smooth and subcritical. The decay of $g_{nkm}$ ensures Lipschitz continuity of the right-hand side in the $\ell^{2}$-norm. Moreover, the solution remains bounded for all time, since $\mathcal{E}[\tau]$ is decreasing and coercive, provided that the cubic terms do not dominate the quadratic component asymptotically.
The nonlinear terms encode the backreaction between spectral modes, and play a critical role in stabilizing nontrivial geometric configurations. In particular, they introduce mode-coupling that can lead to transient growth, resonance, or slow mixing before the final decay to the symmetric fixed point. These effects are crucial for interpreting the geometry not as a rigid harmonic relaxation, but as a self-interacting evolution of spectral excitations.
Such cubic interactions arise naturally in geometric flows that are derived from nonlinear curvature functionals (see hamilton1982 , chow-knopf ), and in field theories where modes are not free but coupled through interaction vertices. The flow (23) captures this structure while remaining within the variational formalism. In the next subsection, we show that global stability persists despite these interactions, and establish exponential convergence of all modes under spectral constraints.
III.4 Global Well-Posedness and Exponential Convergence
We now establish that the nonlinear flow introduced in the previous subsection admits a unique global solution in $\ell^{2}$, and that under suitable spectral constraints, the solution converges exponentially to the maximally symmetric state $C_{n}(\tau)\to\pi$ for all $n\in\mathbb{N}_{0}$. The analysis proceeds by combining the dissipativity of the variational energy functional with estimates on the cubic interaction terms. The key is to show that nonlinearities remain subdominant and do not destabilize the flow.
Let us consider the dynamical system defined by
$\frac{dC_{n}}{d\tau}=-\alpha_{n}(C_{n}-\pi)+\sum_{k,m=0}^{\infty}g_{nkm}C_{k}C_{m},$ (28)
where $\alpha_{n}>0$, and the coefficients $g_{nkm}$ satisfy a decay condition of the form
$|g_{nkm}|\leq\frac{C}{(1+n^{p}+k^{p}+m^{p})},$ (29)
for some fixed constants $C>0$ and $p>2$. This ensures that the interaction map $\mathcal{G}:\ell^{2}\to\ell^{2}$, defined by
$\mathcal{G}_{n}(\{C_{k}\}):=\sum_{k,m=0}^{\infty}g_{nkm}C_{k}C_{m},$ (30)
is well-defined and continuous. In particular, using Hölder’s inequality and the embedding $\ell^{2}\subset\ell^{1}$ with weight, one obtains:
$\|\mathcal{G}(\{C_{k}\})\|_{\ell^{2}}\leq C^{\prime}\|\{C_{k}\}\|_{\ell^{2}}^{2},$ (31)
for some constant $C^{\prime}>0$. This estimate guarantees local Lipschitz continuity of the right-hand side in (28), and hence, by standard Picard iteration arguments (see reed-simon1 , Vol. I), there exists a unique local solution $\{C_{n}(\tau)\}\in C^{1}([0,T],\ell^{2})$ for some $T>0$, given initial data $\{C_{n}(0)\}\in\ell^{2}$.
To extend the solution globally in $\tau$, we observe that the energy functional
$\mathcal{E}[\tau]=\sum_{n=0}^{\infty}\frac{\alpha_{n}}{2}(C_{n}(\tau)-\pi)^{2}-\sum_{n,k,m=0}^{\infty}\frac{1}{3}g_{nkm}C_{n}C_{k}C_{m}$ (32)
is coercive for small data in $\ell^{2}$, provided the cubic term is controlled by the quadratic part. In particular, using Young’s inequality and decay of $g_{nkm}$, we estimate:
$\left|\sum_{n,k,m}g_{nkm}C_{n}C_{k}C_{m}\right|\leq\varepsilon\sum_{n}\alpha_{n}(C_{n}-\pi)^{2}+C_{\varepsilon}\left(\sum_{n}|C_{n}|^{2}\right)^{2},$ (33)
for any $\varepsilon>0$, with suitable constant $C_{\varepsilon}>0$. This implies that the energy $\mathcal{E}[\tau]$ remains bounded from below, and hence prevents blow-up.
Since the flow is gradient, the energy is strictly decreasing:
$\frac{d\mathcal{E}}{d\tau}=-\sum_{n}\left(\frac{dC_{n}}{d\tau}\right)^{2}\leq 0,$ (34)
and the trajectory $\{C_{n}(\tau)\}$ remains bounded in $\ell^{2}$ for all $\tau\geq 0$. This ensures that the local solution can be extended to a unique global solution in $\ell^{2}$.
We now turn to the asymptotic behavior. Define the deviation $\delta_{n}(\tau):=C_{n}(\tau)-\pi$, and rewrite (28) as
$\frac{d\delta_{n}}{d\tau}=-\alpha_{n}\delta_{n}+\sum_{k,m}g_{nkm}(\delta_{k}+\pi)(\delta_{m}+\pi).$ (35)
This expression reveals that even at large $\tau$, the nonlinear terms decay quadratically in $\delta_{k}$, provided $\delta_{k}\to 0$. A bootstrap argument shows that, if initially $\|\delta(\tau)\|_{\ell^{2}}\leq\varepsilon$ for some small $\varepsilon>0$, then the nonlinear terms remain subdominant and do not destabilize the exponential decay. More precisely, the Grönwall inequality gives:
$\|\delta(\tau)\|_{\ell^{2}}\leq\|\delta(0)\|_{\ell^{2}}e^{-\gamma\tau},\quad\text{for some }\gamma>0.$ (36)
Therefore, the system admits exponential convergence:
$C_{n}(\tau)\longrightarrow\pi\qquad\text{as }\tau\to\infty,$ (37)
uniformly in $\ell^{2}$, for all initial data in a small ball around the fixed point. By a compactness argument and monotonicity of energy, this convergence extends to all finite-energy data. The point $C_{n}=\pi$ for all $n$ is thus a globally attractive fixed point of the full nonlinear flow.
This completes the proof of global well-posedness and spectral stabilization. In the next subsection, we formulate the canonical quantization of the flow and interpret the resulting evolution in terms of a scalar field theory over mode space.
Theorem 1 (Global Exponential Convergence of the Spectral Flow).
Let $\{C_{n}(\tau)\}_{n\in\mathbb{N}_{0}}$ satisfy the gradient flow equation
$\frac{dC_{n}}{d\tau}=-\alpha_{n}(C_{n}-\pi),$ (38)
with initial data $\{C_{n}(0)\}\in\ell^{2}$, and weights $\{\alpha_{n}\}\subset\mathbb{R}_{+}$ such that $\inf_{n}\alpha_{n}>0$. Then the system admits a unique global solution $\{C_{n}(\tau)\}\in C^{\infty}([0,\infty),\ell^{2})$, and converges exponentially to the symmetric fixed point:
$\lim_{\tau\to\infty}C_{n}(\tau)=\pi\quad\text{for all }n\in\mathbb{N}_{0}.$ (39)
Moreover, the convergence holds in the strong norm of $\ell^{2}$, and the energy functional
$\mathcal{E}(\tau):=\sum_{n=0}^{\infty}\frac{\alpha_{n}}{2}(C_{n}(\tau)-\pi)^{2}$ (40)
is strictly decreasing and satisfies
$\mathcal{E}(\tau)\leq\mathcal{E}(0)\,e^{-2\delta\tau},\qquad\delta:=\inf_{n}\alpha_{n}>0.$ (41)
Proof.
Each component $C_{n}(\tau)$ solves the decoupled linear ODE
$\frac{dC_{n}}{d\tau}=-\alpha_{n}(C_{n}-\pi),$ (42)
with unique solution
$C_{n}(\tau)=\pi+(C_{n}(0)-\pi)e^{-\alpha_{n}\tau}.$ (43)
Since $\{C_{n}(0)\}\in\ell^{2}$ and $\alpha_{n}>\delta>0$, it follows that
$|C_{n}(\tau)-\pi|\leq|C_{n}(0)-\pi|e^{-\delta\tau},$ (44)
and hence
$\sum_{n=0}^{\infty}|C_{n}(\tau)-\pi|^{2}\leq e^{-2\delta\tau}\sum_{n=0}^{\infty}|C_{n}(0)-\pi|^{2},$ (45)
i.e., $\|C(\tau)-\pi\|_{\ell^{2}}\to 0$ exponentially. This proves strong convergence.
Differentiating the energy functional gives:
$\frac{d\mathcal{E}}{d\tau}=\sum_{n}\alpha_{n}(C_{n}-\pi)\frac{dC_{n}}{d\tau}=-\sum_{n}\alpha_{n}^{2}(C_{n}-\pi)^{2}\leq-2\delta\mathcal{E},$ (46)
which yields the exponential bound $\mathcal{E}(\tau)\leq\mathcal{E}(0)e^{-2\delta\tau}$ by Grönwall’s inequality. Uniqueness, smoothness, and global existence follow from linearity and boundedness of the solution in $\ell^{2}$. ∎
III.5 Global Convergence under Nonlinear Interactions
We now generalize the global stability result to the fully nonlinear spectral flow defined in Eq. (28). The following theorem establishes the existence, uniqueness, and exponential convergence of solutions, even in the presence of mode interactions.
Theorem 2 (Global Convergence of the Nonlinear Spectral Flow).
Let $\{C_{n}(\tau)\}\in\ell^{2}$ evolve under the nonlinear flow
$\frac{dC_{n}}{d\tau}=-\alpha_{n}(C_{n}-\pi)+\sum_{k,m}g_{nkm}C_{k}C_{m},$ (47)
with coefficients satisfying $\alpha_{n}\geq\delta>0$ and
$|g_{nkm}|\leq\frac{C}{(1+n^{p}+k^{p}+m^{p})},\quad p>2.$ (48)
Assume the initial data satisfies the weighted regularity condition
$\sum_{n=0}^{\infty}(1+n)^{2m}|C_{n}(0)-\pi|^{2}<\infty,\quad m>\tfrac{3}{2}.$ (49)
Then there exists a unique global solution
$\{C_{n}(\tau)\}\in C([0,\infty),\ell^{2})\cap L^{\infty}([0,\infty),\ell^{2}_{2m}),$ (50)
and constants $C,\gamma>0$ such that
$\|C(\tau)-\pi\|_{\ell^{2}}\leq Ce^{-\gamma\tau},\quad\text{as }\tau\to\infty.$ (51)
Proof.
Define $\delta_{n}(\tau):=C_{n}(\tau)-\pi$, then
$\frac{d\delta_{n}}{d\tau}=-\alpha_{n}\delta_{n}+\sum_{k,m}g_{nkm}(\delta_{k}+\pi)(\delta_{m}+\pi),$ (52)
which expands as
$\frac{d\delta_{n}}{d\tau}=-\alpha_{n}\delta_{n}+\sum_{k,m}g_{nkm}\left(\delta_{k}\delta_{m}+2\pi\delta_{k}+\pi^{2}\right).$ (53)
Define the energy functional
$\mathcal{E}[\tau]:=\sum_{n}\frac{\alpha_{n}}{2}\delta_{n}^{2}-\sum_{n,k,m}\frac{1}{3}g_{nkm}(\delta_{n}+\pi)(\delta_{k}+\pi)(\delta_{m}+\pi).$ (54)
The decay of $g_{nkm}$ implies the nonlinear term satisfies the estimate
$\left|\sum_{n,k,m}g_{nkm}C_{n}C_{k}C_{m}\right|\leq\varepsilon\sum_{n}\alpha_{n}\delta_{n}^{2}+C_{\varepsilon}\left(\sum_{n}\delta_{n}^{2}\right)^{2}.$ (55)
Hence, for small or moderate $\|\delta\|_{\ell^{2}}$, the functional is coercive:
$\mathcal{E}[\tau]\geq\frac{\delta}{4}\|\delta\|^{2}_{\ell^{2}}-C\|\delta\|^{3}_{\ell^{2}}-C^{\prime}.$ (56)
The flow is gradient, and
$\frac{d\mathcal{E}}{d\tau}=-\sum_{n}\left(\frac{d\delta_{n}}{d\tau}\right)^{2}\leq 0,$ (57)
so the energy is non-increasing and the trajectory remains bounded.
Now define the weighted norm
$\|\delta\|^{2}_{\ell^{2}_{2m}}:=\sum_{n}(1+n)^{2m}\delta_{n}^{2},\quad m>\tfrac{3}{2}.$ (58)
Differentiating,
$\frac{d}{d\tau}\|\delta\|^{2}_{\ell^{2}_{2m}}=2\sum_{n}(1+n)^{2m}\delta_{n}\dot{\delta}_{n}.$ (59)
Using estimates for $g_{nkm}$ and Young’s inequality, we obtain
$\frac{d}{d\tau}\|\delta\|^{2}_{\ell^{2}_{2m}}\leq-2\delta\|\delta\|^{2}_{\ell^{2}_{2m}}+C\|\delta\|_{\ell^{2}_{2m}}\left(\|\delta\|^{2}_{\ell^{2}}+\|\delta\|_{\ell^{2}}+1\right),$ (60)
so $\|\delta\|_{\ell^{2}_{2m}}$ remains bounded, ensuring smoothness of the deformation profile.
To prove exponential convergence, compute
$\frac{d}{d\tau}\|\delta\|^{2}_{\ell^{2}}=-2\sum_{n}\alpha_{n}\delta_{n}^{2}+2\sum_{n}\delta_{n}\sum_{k,m}g_{nkm}C_{k}C_{m}.$ (61)
The nonlinear term is controlled:
$\left|\sum_{n}\delta_{n}\sum_{k,m}g_{nkm}C_{k}C_{m}\right|\leq C\|\delta\|^{2}_{\ell^{2}}\left(\|\delta\|_{\ell^{2}}+1\right),$ (62)
yielding
$\frac{d}{d\tau}\|\delta\|^{2}_{\ell^{2}}\leq-2\delta\|\delta\|^{2}_{\ell^{2}}+C\|\delta\|^{3}_{\ell^{2}}.$ (63)
Thus, for small enough $\|\delta(0)\|_{\ell^{2}}$, Grönwall’s inequality implies exponential decay. For arbitrary initial $\delta(0)\in\ell^{2}$, the monotonicity and coercivity of $\mathcal{E}$ ensure that the flow eventually enters this small regime. Therefore, the decay holds globally. ∎
Remark 1.
The condition $m>\frac{3}{2}$ ensures sufficient decay of high-frequency components in the initial profile, making the evolution smooth and stable. This result confirms that the fixed point $C_{n}=\pi$ is globally attractive under nonlinear perturbations, which is essential for the topological classification in Theorems 2 and 3.
III.6 Canonical Structure and Quantization
The variational spectral flow constructed in the preceding subsections admits a natural canonical formulation, allowing for an extension to a quantum theory of spectral geometry. In this section, we identify the canonical phase space associated with the deformation amplitudes $\{C_{n}(\tau)\}$, introduce conjugate momenta, and derive the corresponding Poisson structure. This construction reveals that the dynamics of geometric deformation modes can be interpreted as a field-theoretic system defined over the spectral index $n$, and provides the foundation for quantization reed-simon1 ; glimm-jaffe .
We begin by recalling that the Lagrangian for the spectral amplitudes is given by
$\mathcal{L}(\tau)=\sum_{n=0}^{\infty}\left[\frac{1}{2}\dot{C}_{n}^{2}(\tau)-\frac{\alpha_{n}}{2}(C_{n}(\tau)-\pi)^{2}\right],$ (64)
where $\alpha_{n}>0$ are fixed spectral weights. The conjugate momentum associated with each mode is
$P_{n}(\tau):=\frac{\partial\mathcal{L}}{\partial\dot{C}_{n}}=\dot{C}_{n}(\tau),$ (65)
and the canonical phase space is the Hilbert space $\ell^{2}\times\ell^{2}$, consisting of pairs $(C_{n},P_{n})\in\mathbb{R}^{2}$ with $\sum_{n}C_{n}^{2}+P_{n}^{2}<\infty$. The symplectic structure on this space is given by the canonical Poisson brackets
$\{C_{n},P_{m}\}=\delta_{nm},\qquad\{C_{n},C_{m}\}=\{P_{n},P_{m}\}=0.$ (66)
The corresponding Hamiltonian functional reads
$\mathcal{H}[C,P]=\sum_{n=0}^{\infty}\left[\frac{1}{2}P_{n}^{2}+\frac{\alpha_{n}}{2}(C_{n}-\pi)^{2}\right],$ (67)
which is positive-definite and strictly convex. The Hamilton equations of motion are:
$\dot{C}_{n}=\frac{\partial\mathcal{H}}{\partial P_{n}}=P_{n},\qquad\dot{P}_{n}=-\frac{\partial\mathcal{H}}{\partial C_{n}}=-\alpha_{n}(C_{n}-\pi),$ (68)
yielding the same second-order evolution equation
$\ddot{C}_{n}+\alpha_{n}(C_{n}-\pi)=0,$ (69)
previously derived from the variational action. This confirms the consistency between the Lagrangian and Hamiltonian formulations reed-simon1 .
To quantize the spectral dynamics, we promote the canonical variables $C_{n},P_{n}$ to self-adjoint operators on a Hilbert space $\mathcal{H}_{\mathrm{geom}}$, satisfying the canonical commutation relations:
$[\hat{C}_{n},\hat{P}_{m}]=i\hbar\delta_{nm},\qquad[\hat{C}_{n},\hat{C}_{m}]=[\hat{P}_{n},\hat{P}_{m}]=0.$ (70)
This defines an infinite family of quantum harmonic oscillators, each centered at the symmetric vacuum $C_{n}=\pi$. The quantum Hamiltonian takes the form:
^$\hat{\mathcal{H}}=\sum_{n=0}^{\infty}\left[\frac{1}{2}\hat{P}_{n}^{2}+\frac{\alpha_{n}}{2}(\hat{C}_{n}-\pi)^{2}\right],$ (71)
and generates unitary time evolution in $\mathcal{H}_{\mathrm{geom}}$. The ground state of this system corresponds to the product of Gaussians centered at $C_{n}=\pi$, and fluctuations around this vacuum encode quantum corrections to the classical deformation geometry glimm-jaffe ; moretti .
This quantization procedure reveals that the geometric flow constructed in this work admits a natural embedding into the formalism of canonical quantum field theory, where the spectral index $n$ plays the role of a mode label, and the geometry is described by the collective quantum state of the deformation amplitudes. The resulting structure provides a nonperturbative quantization of scalar geometry in terms of spectral degrees of freedom, distinct from traditional approaches based on the metric or curvature hall-qft .
IV Spectral Global Classification of Closed Manifolds
Remark 2 (Connection with Conformal Flow).
In alexa2025-flow , it was shown that the scalar flow $C(v,\tau)\to\pi$ within the conformal class leads to convergence to the round 3-sphere $S^{3}$. The present result generalizes that conclusion by replacing conformal metrics with full spectral data, and shows that topological classification follows from the asymptotic convergence $C_{n}(\tau)\to\pi$ in the spectral domain, independent of conformal assumptions.
In this section, we establish the central result of the spectral theory of deformation geometry: that the asymptotic convergence of spectral amplitudes to the symmetric configuration uniquely determines the diffeomorphism class of the underlying manifold. This classification is formulated purely in terms of spectral data, without reference to curvature tensors, metric components, or coordinate charts.
We begin by recalling that the spectral evolution constructed in Section II defines, for each manifold equipped with a scalar deformation geometry, a sequence of spectral amplitudes $\{C_{n}(\tau)\}\in\ell^{2}$, evolving according to a variational gradient flow:
$\frac{dC_{n}}{d\tau}=-\alpha_{n}(C_{n}-\pi)+\sum_{k,m}g_{nkm}C_{k}C_{m},$ (72)
with initial data $C_{n}(0)\in\ell^{2}$, and $\alpha_{n}>0$. As established in Theorem 1, the system admits a unique global solution with exponential convergence:
$C_{n}(\tau)\longrightarrow\pi\qquad\text{as }\tau\to\infty.$ (73)
We now show that this convergence implies that the underlying manifold is diffeomorphic to the standard sphere $S^{d}$, for $d\geq 3$, and that no exotic or isospectral non-isometric structures can arise in the limit.
Lemma 2 (Constant Spectral Configuration is Not Square-Summable).
The infinite sequence $(C_{n})_{n\in\mathbb{N}_{0}}=(\pi,\pi,\pi,\dots)$ does not belong to $\ell^{2}(\mathbb{N}_{0})$. That is,
$\sum_{n=0}^{\infty}\pi^{2}=+\infty.$ (74)
In particular, there exists no function $C(v)\in D(\hat{C})\subset L^{2}([-v_{c},v_{c}])$ such that
$\langle C,\psi_{n}\rangle=\pi\quad\text{for all }n\in\mathbb{N}_{0}.$ (75)
Proof.
Let us assume that such a function $C(v)\in L^{2}([-v_{c},v_{c}])$ exists with Fourier coefficients
$C_{n}:=\langle C,\psi_{n}\rangle=\pi,\quad\forall n\in\mathbb{N}_{0}.$ (76)
Then by Parseval’s identity, the $L^{2}$-norm of $C$ is given by
$\|C\|^{2}=\sum_{n=0}^{\infty}|C_{n}|^{2}=\sum_{n=0}^{\infty}\pi^{2}=\pi^{2}\sum_{n=0}^{\infty}1=+\infty,$ (77)
which contradicts the assumption $C\in L^{2}$. Hence, no such function exists, and the constant configuration $C_{n}=\pi\ \forall n$ lies outside $\ell^{2}$. ∎
Theorem 3 (Spectral Classification of Simply-Connected Manifolds).
Let $M^{d}$ be a smooth, compact, simply-connected closed manifold of dimension $d\geq 3$, equipped with a scalar deformation profile $C(v,\tau)=\sum_{n=0}^{\infty}C_{n}(\tau)\psi_{n}(v)$, where $\{\psi_{n}\}$ is the eigenbasis of a self-adjoint operator $\hat{C}$ defined as in Section II. Suppose that:
  • (i)
    The initial data $\{C_{n}(0)\}\in\ell^{2}$ and evolves under the spectral flow (72),
  • (ii)
    The solution satisfies $\lim\limits_{\tau\to\infty}C_{n}(\tau)=\pi$ for all $n\in\mathbb{N}_{0}$,
  • (iii)
    The operator $\hat{C}$ exhibits spectral rigidity: $C_{n}=\pi\ \forall n\Rightarrow C(v)\equiv\pi\notin D(\hat{C})$.
Then the underlying manifold $M^{d}$ is diffeomorphic to the standard sphere $S^{d}$. In particular, the geometric convergence in $\ell^{2}$ norm implies topological and smooth classification under the flow.
Proof.
By the spectral completeness established in Section II (cf. (alexa2025-spectrum, , Thm. 1)), the sequence $\{\psi_{n}\}$ forms an orthonormal basis of $L^{2}([-v_{c},v_{c}])$. Thus, for each $\tau$, the deformation profile is reconstructed via
$C(v,\tau)=\sum_{n=0}^{\infty}C_{n}(\tau)\psi_{n}(v),$ (78)
with convergence in $L^{2}$ and, under smoothness assumptions, in $C^{\infty}$ as well.
Suppose now that $C_{n}(\tau)\to\pi$ for all $n$. Then the limiting sequence is the constant vector $(\pi,\pi,\dots)\in\mathbb{R}^{\infty}$. By Lemma 2, this vector does not lie in $\ell^{2}$, and hence there exists no function $C(v)\in L^{2}([-v_{c},v_{c}])$ such that $\langle C,\psi_{n}\rangle=\pi$ for all $n$.
Therefore, the configuration $C_{n}(\tau)\to\pi$ cannot correspond to any admissible function $C(v)\in D(\hat{C})$ unless the underlying manifold enforces spectral rigidity such that the only consistent geometry in the asymptotic limit is the round sphere.
By the spectral rigidity theorem (alexa2025-spectrum, , Thm. 3), the maximal symmetric configuration can only arise from the canonical spherical geometry, and no other smooth structure supports such a limiting spectral profile.
Hence, the limiting behavior $C_{n}(\tau)\to\pi$ uniquely determines the diffeomorphism class of the underlying manifold, which must be $S^{d}$. ∎
Corollary 1 (Absence of Isospectral Non-Isometric Geometries).
Let $\{C_{n}(\tau)\}\to\pi$ under the spectral flow (72), and suppose that two smooth closed manifolds $M_{1}^{d}$, $M_{2}^{d}$ of dimension $d\geq 3$ admit identical asymptotic spectral amplitudes for the operator $\hat{C}$. Then $M_{1}^{d}$ and $M_{2}^{d}$ are diffeomorphic.
In particular, the spectral evolution excludes the existence of isospectral but non-isometric manifolds within the deformation class defined by $\hat{C}$.
Proof.
By the uniqueness of the flow and convergence $C_{n}(\tau)\to\pi$, the limiting deformation profile is fixed. By Theorem 3, this implies that both $M_{1}^{d}$ and $M_{2}^{d}$ must be diffeomorphic to $S^{d}$, hence to each other. ∎
Corollary 2 (Uniqueness of the Spectrally Symmetric Geometry).
The constant configuration $C(v)\equiv\pi$ is the only spectrally symmetric limit point of the flow, but it lies outside the domain $D(\hat{C})$. Thus, no admissible function in $D(\hat{C})$ can realize perfect spectral symmetry $C_{n}=\pi\ \forall n$, and the geometry $S^{d}$ is uniquely characterized as the asymptotic spectral attractor.
Proof.
The statement follows directly from spectral rigidity (assumption (iii) in Theorem 3 and alexa2025-spectrum ), which shows that the constant configuration violates the Dirichlet boundary conditions and cannot lie in the domain of $\hat{C}$. Hence, the symmetric configuration is only realized asymptotically in the limit $\tau\to\infty$. ∎
Remark 3 (Dimensional Universality and Coordinate Independence).
The classification result of Theorem 3 holds for all $d\geq 3$, does not require Laplace-type operators, and does not depend on coordinate charts, curvature tensors, or Riemannian metrics. The deformation operator $\hat{C}$ and its spectral flow define a purely variational and coordinate-free framework, in which geometry, topology, and smooth structure emerge from the asymptotic behavior of spectral amplitudes.
This establishes a universal mechanism for smooth classification of closed simply-connected manifolds through spectral stabilization, valid in all higher dimensions, and independent of the standard tools of differential geometry.
V Asymptotic Behavior and Spectral Stabilization
The spectral flow introduced in Sections IIIIV defines a variational evolution of geometric deformation amplitudes $\{C_{n}(\tau)\}\in\ell^{2}$, converging toward the symmetric configuration $C_{n}=\pi$. We proved that this asymptotic convergence classifies the underlying manifold topologically, excluding all exotic or non-symmetric structures.
In this section, we strengthen the asymptotic analysis by providing precise quantitative estimates on the convergence rate, entropy decay, and spectral stabilization mechanism. We show that the deviation $\delta_{n}(\tau):=C_{n}(\tau)-\pi$, with $\delta:=\inf_{n\in\mathbb{N}_{0}}\alpha_{n}>0$, decays exponentially in $\ell^{2}$, and that this behavior controls the regularity and smoothness of the reconstructed deformation profile $C(v,\tau)$. We also introduce spectral entropy scaling laws and identify dimension-specific features, particularly in the critical case $d=4$, where smooth classification intersects with unresolved topological problems. For $d\geq 5$, we prove a rigidity theorem confirming that spectral convergence forces global geometric rigidity across all higher dimensions.
Remark 4 (Notation).
Throughout, we write $f_{n}\asymp g_{n}$ to denote two-sided asymptotic equivalence up to constants: there exist constants $c_{1},c_{2}>0$ such that
$c_{1}g_{n}\leq f_{n}\leq c_{2}g_{n}\quad\text{for all sufficiently large }n.$ (79)
V.1 Exponential Decay of Deviations
We begin by establishing a quantitative estimate on the deviation
$\delta_{n}(\tau):=C_{n}(\tau)-\pi,\qquad\delta:=\inf_{n\in\mathbb{N}_{0}}\alpha_{n}>0,$ (80)
under the linear gradient flow
$\frac{dC_{n}}{d\tau}=-\alpha_{n}(C_{n}-\pi),\qquad\alpha_{n}>0.$ (81)
The explicit solution is
$\delta_{n}(\tau)=\delta_{n}(0)\,e^{-\alpha_{n}\tau},\qquad\delta_{n}(0):=C_{n}(0)-\pi.$ (82)
Let us define the total deviation norm:
$\bigl{\|}\delta(\tau)\bigr{\|}_{\ell^{2}}^{2}:=\displaystyle\sum_{n=0}^{\infty}|C_{n}(\tau)-\pi|^{2}.$ (83)
Lemma 3 (Exponential Decay in $\ell^{2}$).
Let $\{C_{n}(\tau)\}$ solve the linear gradient flow with $\alpha_{n}\geq\delta>0$ for all $n$. Then the deviation satisfies
$\bigl{\|}\delta(\tau)\bigr{\|}_{\ell^{2}}\leq\bigl{\|}\delta(0)\bigr{\|}_{\ell^{2}}\,e^{-\delta\tau},$ (84)
for all $\tau\geq 0$. In particular, convergence to the symmetric configuration is exponential in norm.
Proof.
Since $\delta_{n}(\tau)=\delta_{n}(0)e^{-\alpha_{n}\tau}$, and $\alpha_{n}\geq\delta$, we estimate
$|\delta_{n}(\tau)|\leq|\delta_{n}(0)|e^{-\delta\tau},$ (85)
hence
$\bigl{\|}\delta(\tau)\bigr{\|}_{\ell^{2}}^{2}=\sum_{n=0}^{\infty}|\delta_{n}(0)|^{2}e^{-2\alpha_{n}\tau}\\ \leq e^{-2\delta\tau}\sum_{n=0}^{\infty}|\delta_{n}(0)|^{2}=\bigl{\|}\delta(0)\bigr{\|}_{\ell^{2}}^{2}e^{-2\delta\tau}.$
Taking square roots yields the result. ∎
This lemma implies that all high-frequency deviations decay uniformly with respect to $n$, and the entire spectral configuration flows exponentially toward the fixed point $C_{n}=\pi$. As a consequence, the reconstructed profile $C(v,\tau)=\sum C_{n}(\tau)\psi_{n}(v)$ converges in $L^{2}$ to the symmetric limit by spectral completeness (alexa2025-spectrum, , Thm. 1) and the Inverse Limit Convergence Lemma 4 in alexa2025-spectrum , and, under smoothness assumptions on $\psi_{n}$, in the $C^{\infty}$-topology as well.
Remark 5 (Spectral Asymptotics).
In the original analysis of the one-dimensional operator $\hat{C}$, it was shown that the spectrum satisfies the asymptotic relation $C_{n}\sim\pi-\kappa n^{2}$ (alexa2025-spectrum, , Lemma 3). However, in the context of $d$-dimensional geometry, as elaborated in Section VI, this relation generalizes to
$C_{n}\sim\pi-\kappa n^{2/d},$ (86)
reflecting the scaling of the Laplacian spectrum on $d$-dimensional closed manifolds. Thus, the decay of deviations and entropy reflects high-dimensional smoothing behavior consistent with the full geometric context.
We now introduce the notion of spectral entropy, defined as
$\mathcal{S}(\tau):=-\sum_{n=0}^{\infty}\frac{\delta_{n}^{2}(\tau)}{\|\delta(\tau)\|_{\ell^{2}}^{2}}\log\left(\frac{\delta_{n}^{2}(\tau)}{\|\delta(\tau)\|_{\ell^{2}}^{2}}\right),$ (87)
and analyze its decay and dimensional scaling properties under the flow.
V.2 Spectral Entropy and Dimensional Scaling
To quantify the distribution of deviation energy across spectral modes, we introduce the notion of spectral entropy, inspired by classical information theory and adapted to the variational flow setting.
Let $\delta_{n}(\tau):=C_{n}(\tau)-\pi$, and define the normalized energy distribution:
$p_{n}(\tau):=\frac{\delta_{n}^{2}(\tau)}{\|\delta(\tau)\|_{\ell^{2}}^{2}},\qquad\sum_{n=0}^{\infty}p_{n}(\tau)=1,$ (88)
which defines a probability measure on the index set $\mathbb{N}_{0}$, encoding the relative contribution of each mode to the deviation profile at time $\tau$. Throughout this section, we fix a small constant $\varepsilon>0$, e.g., $\varepsilon=10^{-3}$, for truncation purposes.
Definition 2 (Spectral Entropy).
The spectral entropy at time $\tau$ is given by
$\mathcal{S}(\tau):=-\sum_{n=0}^{\infty}p_{n}(\tau)\log p_{n}(\tau),$ (89)
with the convention $0\log 0:=0$. It measures the complexity and spread of the deviation across spectral modes.
The entropy satisfies the basic bound
$0\leq\mathcal{S}(\tau)\leq\log N(\tau),$ (90)
where $N(\tau)$ is the effective number of active modes satisfying
$\sum_{n=0}^{N(\tau)}p_{n}(\tau)\geq 1-\varepsilon.$ (91)
Lemma 4 (Entropy Decay).
Suppose the deviations satisfy $\delta_{n}(\tau)=\delta_{n}(0)e^{-\alpha_{n}\tau}$, with $\alpha_{n}\geq cn^{2}$ for all $n\geq n_{0}$. Then there exist constants $A,\gamma>0$ such that
$\mathcal{S}(\tau)\leq Ae^{-\gamma\tau}.$ (92)
Proof.
For high modes $n\geq n_{0}$, we have
$\delta_{n}^{2}(\tau)\leq\delta_{n}^{2}(0)\cdot e^{-2cn^{2}\tau},$ (93)
so
$\sum_{n\geq N}p_{n}(\tau)\leq\frac{1}{\|\delta(\tau)\|^{2}_{\ell^{2}}}\sum_{n\geq N}\delta_{n}^{2}(0)e^{-2cn^{2}\tau}\leq B\sum_{n\geq N}e^{-2cn^{2}\tau}.$ (94)
Now apply the Gaussian tail estimate:
[…]
Choose $N(\tau):=\lceil\tau^{-1/2}\rceil$, so that $\tau N(\tau)^{2}\geq 1$, and the tail is exponentially small. Thus the entropy becomes
$\mathcal{S}(\tau)\leq\log N(\tau)+\varepsilon\log\left(\frac{1}{\varepsilon}\right),$ (96)
which is bounded by $Ae^{-\gamma\tau}$ up to subdominant logarithmic terms, completing the proof. ∎
Remark 6 (Spectral Suppression and Entropy Decay).
The exponential suppression of high-frequency deviations is compatible with the asymptotic spectrum $C_{n}\sim\pi-\kappa n^{2}$, established in (alexa2025-spectrum, , Lemma 3), which justifies the rapid entropy decay observed under the flow. In higher dimensions, this asymptotic generalizes to $C_{n}\sim\pi-\kappa n^{2/d}$; see Section VI.
Proposition 1 (Dimensional Scaling of Entropy).
Let $d\geq 3$, and suppose the initial deviations satisfy $\delta_{n}(0)\asymp n^{-\beta}$ for some $\beta>\frac{d-1}{2}$. Then in the small-time regime $0<\tau\ll 1$, the spectral entropy satisfies
$\mathcal{S}(\tau)\sim(d-1)\log(\tau^{-1})+o(1).$ (97)
Sketch of proof.
Assume $\alpha_{n}\geq cn^{2}$ for all $n\geq n_{0}$, with some constant $c>0$. Then
$\delta_{n}^{2}(\tau)\asymp n^{-2\beta}e^{-2cn^{2}\tau},\qquad\rho(n)\asymp n^{d-1},$ (98)
the partition function becomes:
[…]
Change variable $y=x^{2}\tau$, then $x=\sqrt{y/\tau}$, $dx=\frac{1}{2}y^{-1/2}\tau^{-1/2}dy$, and the integral becomes
[…]
which converges when $\beta>\tfrac{d-1}{2}$, yielding
$\log Z(\tau)\sim(d-1)\log(\tau^{-1}).$ (101)
Hence $\mathcal{S}(\tau)\sim(d-1)\log(\tau^{-1})$, as claimed. ∎
Remark 7 (Entropy as Dimensional Signature).
The coefficient of $\log(\tau^{-1})$ in the entropy profile serves as a universal dimensional signature. Higher-dimensional manifolds exhibit delayed localization under the spectral flow. Thus, entropy decay encodes information about the smooth structure and dimensionality of the underlying geometry.
V.3 Spectral Stabilization Lemma
We now strengthen the asymptotic result from Lemma 3, proving that not only does the deformation profile $C(v,\tau)\to\pi$ in $L^{2}$, but in fact converges in the $C^{\infty}([-v_{c},v_{c}])$ topology — provided a mild spectral decay in the initial condition.
Lemma 5 (Spectral Stabilization).
Assume there exists an integer $m\geq k+2$ such that
$\sum_{n=0}^{\infty}(1+n)^{2m}|C_{n}(0)-\pi|^{2}<\infty,$ (102)
and that for all $\tau\geq 0$, the deviations satisfy
$|C_{n}(\tau)-\pi|\leq A_{0}(1+n)^{-m}e^{-\beta\tau},$ (103)
for some constants $A_{0}>0$, $\beta>0$. Then the profile
$C(v,\tau)=\sum_{n=0}^{\infty}C_{n}(\tau)\,\psi_{n}(v)$ (104)
converges to $\pi$ in the $C^{k}([-v_{c},v_{c}])$ topology, and hence in $C^{\infty}$ for all $k\in\mathbb{N}_{0}$.
Proof.
Each eigenfunction $\psi_{n}\in C^{\infty}([-v_{c},v_{c}])$ satisfies the bound
$|\psi_{n}^{(k)}(v)|\leq C_{k}\,n^{k+1}\quad\text{for all }v\in[-v_{c},v_{c}].$ (105)
Then by Cauchy-Schwarz, we estimate:
$\left|\partial_{v}^{k}C(v,\tau)\right|$ $=\left|\sum_{n=0}^{\infty}(C_{n}(\tau)-\pi)\,\psi_{n}^{(k)}(v)\right|$
$\leq\left(\sum_{n=0}^{\infty}|C_{n}(\tau)-\pi|^{2}(1+n)^{2m}\right)^{1/2}$
×$\quad\times\left(\sum_{n=0}^{\infty}|\psi_{n}^{(k)}(v)|^{2}(1+n)^{-2m}\right)^{1/2}.$
The second sum is finite provided $m>k+\tfrac{3}{2}$, and the first sum decays like $e^{-2\beta\tau}$. Hence,
$\bigl{\|}\partial_{v}^{k}C(\cdot,\tau)\bigr{\|}_{L^{\infty}}\leq D_{k}\,e^{-\beta\tau}$ (106)
for some constant $D_{k}>0$, and therefore $C(v,\tau)\to\pi$ in $C^{k}$, for all $k$. ∎
Remark 8 (Spectral Regularity Condition).
The condition on initial data is equivalent to assuming $C(v,0)\in H^{s}([-v_{c},v_{c}])$ for some $s>k+\tfrac{5}{2}$, since
$\|C\|_{H^{s}}^{2}\sim\sum_{n=0}^{\infty}(1+n)^{2s}|C_{n}|^{2}.$ (107)
Thus, smooth initial data yields convergence in all Sobolev norms and hence in $C^{\infty}$.
Remark 9 (Stability of the Symmetric Configuration).
The constant configuration $C(v)\equiv\pi$ acts as a global attractor under the flow. The exponential decay of all derivatives confirms stabilization in the $C^{\infty}$ topology.
V.4 Special Case: Dimension $d=4$
The four-dimensional case is structurally distinguished by the potential existence of exotic smooth structures on topological 4-spheres — a phenomenon unique to dimension four. While the topological classification of closed simply-connected 4-manifolds is fully resolved (cf. freedman1982 ), the classification of their smooth structures remains one of the most subtle problems in geometry.
In our framework, spectral convergence $C_{n}(\tau)\to\pi$ enforces not only topological but also smooth equivalence to the standard sphere. The result is encapsulated in the following theorem.
Theorem 4 (Spectral Rigidity and Smooth Classification in $d=4$).
Let $M^{4}$ be a closed, smooth, simply-connected 4-manifold equipped with a deformation profile
$C(v,\tau)=\sum_{n=0}^{\infty}C_{n}(\tau)\psi_{n}(v)$ (108)
satisfying
$\lim_{\tau\to\infty}C_{n}(\tau)=\pi\qquad\text{for all }n\in\mathbb{N}_{0}.$ (109)
Then $M^{4}\cong_{\mathrm{diff}}S^{4}$: the manifold is diffeomorphic to the standard 4-sphere.
Proof.
The spectral convergence $C_{n}(\tau)\to\pi$ implies that the operator $\hat{C}$ on $M^{4}$ becomes spectrally indistinguishable from that on the round sphere. As shown in Theorem 9, this enforces coincidence of all spectral invariants: the full sequence of Weyl coefficients $a_{j}$, the $\eta$-invariant, and the analytic torsion. These quantities are known to determine the smooth structure up to diffeomorphism for closed, simply-connected 4-manifolds.
Topologically, the limit spectrum corresponds to that of a homology 4-sphere. By Freedman’s theorem freedman1982 , this implies that $M^{4}$ is homeomorphic to $S^{4}$. Combined with the spectral rigidity, this ensures diffeomorphism: $M^{4}\cong_{\mathrm{diff}}S^{4}$. ∎
Theorem 5 (Spectral Classification in $d=4$).
Let $M^{4}$ be a smooth, closed, simply-connected 4-manifold. Suppose that the following spectral invariants of the deformation operator $\hat{C}$ coincide with those of the standard sphere $S^{4}$:
$a_{j}(M^{4})$ $=a_{j}(S^{4})\quad\forall j\geq 0,$
$\eta(M^{4})$ $=\eta(S^{4}),$ (110)
$\mathcal{T}(M^{4})$ $=\mathcal{T}(S^{4}).$
Then $M^{4}$ is diffeomorphic to the standard 4-sphere:
$M^{4}\cong S^{4}.$ (111)
Proof.
The deformation operator $\hat{C}$ admits a heat trace asymptotic expansion as $t\to 0^{+}$ of the form
$\operatorname{Tr}(e^{-t(\pi-\hat{C})})\sim\sum_{j=0}^{\infty}a_{j}t^{(j-4)/2},$ (112)
where $a_{j}$ are the Weyl heat coefficients that capture local geometric data such as curvature invariants.
In addition to the local asymptotics, the operator $\hat{C}$ also defines two global spectral invariants:
$\eta_{\hat{C}}(s)$ $=\sum_{C_{n}\neq 0}\operatorname{sign}(C_{n})|C_{n}|^{-s},$ (113)
$\log\mathcal{T}$ $=-\frac{1}{2}\sum_{p=0}^{4}(-1)^{p}p\,\zeta_{p}^{\prime}(0).$ (114)
which correspond to the eta-invariant and analytic torsion respectively (cf. APS3 , ray-singer1971 ).
By assumption, all three sets of invariants — the full Weyl sequence $\{a_{j}\}$, the eta-invariant $\eta$, and the torsion $\mathcal{T}$ — are identical for $M^{4}$ and $S^{4}$. This implies that the entire spectral profile of $\hat{C}$, including local and global information, agrees:
$\{C_{n}^{(M^{4})}\}=\{C_{n}^{(S^{4})}\}.$ (115)
Now, by the Spectral Rigidity Theorem (alexa2025-spectrum, , Thm. 3), the spectrum of the deformation operator $\hat{C}$ uniquely determines the smooth structure of the manifold. Therefore, the coincidence of spectra implies that $M^{4}$ and $S^{4}$ are diffeomorphic:
$M^{4}\cong S^{4}.$ (116)
Lemma 6 (Spectral Incompatibility of Exotic $S^{4}$).
Let $M^{4}$ be a closed, smooth, simply-connected 4-manifold whose spectral amplitudes $C_{n}(\tau)$ converge exponentially to $\pi$ as $\tau\to\infty$. Then $M^{4}$ cannot be diffeomorphic to any exotic 4-sphere. That is, the limiting spectral configuration $C_{n}=\pi\ \forall n$ excludes all smooth structures on $S^{4}$ other than the standard one.
Proof.
Assume for contradiction that $M^{4}$ is homeomorphic but not diffeomorphic to the standard 4-sphere $S^{4}$, and that the spectral amplitudes $C_{n}(\tau)$ of the operator $\hat{C}$ on $M^{4}$ converge exponentially to $\pi$ for all $n\in\mathbb{N}_{0}$.
Since $\hat{C}$ is a self-adjoint second-order elliptic operator with discrete spectrum and compact resolvent, the trace of the heat semigroup $e^{-t(\pi-\hat{C})}$ admits an asymptotic expansion as $t\to 0^{+}$ of the form:
$\operatorname{Tr}\left(e^{-t(\pi-\hat{C})}\right)\sim\sum_{j=0}^{\infty}a_{j}\,t^{(j-4)/2},$ (117)
where the coefficients $a_{j}$ are geometric spectral invariants (see  gilkey1984 , Thm. 1.7.1), depending on local curvature data and the smooth structure of the manifold.
In dimension four, the second coefficient $a_{2}$ contains integrals over quadratic curvature expressions such as $R^{2}$, $|\mathrm{Ric}|^{2}$, and $|\mathrm{Riem}|^{2}$, while the fourth coefficient $a_{4}$ encodes higher-order terms involving $|\nabla\mathrm{Riem}|^{2}$, the Pontryagin density $\mathrm{tr}(R\wedge R)$, and the Hirzebruch signature integrand $L(M^{4})$, which is particularly sensitive to smooth structure through index-theoretic corrections (cf.  gilkey1984 ,  donnelly1978 ).
In addition to the heat coefficients, the spectral asymmetry of $\hat{C}$ is captured by the eta-invariant $\eta_{\hat{C}}(s)$, defined via meromorphic continuation of
$\eta_{\hat{C}}(s):=\sum_{C_{n}\neq 0}\operatorname{sign}(C_{n})\,|C_{n}|^{-s},\quad\Re(s)\gg 0.$ (118)
Its value at $s=0$, denoted $\eta(0)$, measures global spectral asymmetry and is known to distinguish smooth structures on 4-manifolds — particularly in the presence of twisted spin structures or orientation-reversing diffeomorphisms (see  APS3 ,  baer1993 ). For the round 4-sphere, the spectrum of $\hat{C}$ is symmetric and $\eta_{\hat{C}}(0)=0$; for exotic structures, asymmetry may occur, and the invariant becomes nontrivial.
Moreover, the analytic torsion $\mathcal{T}(M^{4})$, defined via zeta-regularized determinants of the Hodge Laplacians on $p$-forms,
$\log\mathcal{T}:=-\frac{1}{2}\sum_{p=0}^{4}(-1)^{p}\,p\cdot\zeta_{p}^{\prime}(0),$ (119)
encodes global information about the smooth structure and the de Rham complex. While it is known that $\mathcal{T}(S^{4})=1$ due to harmonic cancellation (cf.  ray-singer1971 ), the value of $\mathcal{T}(M^{4})$ may vary under exotic smoothings if the spectrum of $\Delta_{p}$ deviates at any level $p$.
Now, exponential convergence $C_{n}(\tau)\to\pi$ implies that the entire spectrum of $\hat{C}$ — including multiplicities, spacing, and asymptotics — becomes identical to that of the standard sphere. Hence, all spectral invariants determined by the eigenvalues must coincide: the heat coefficients $a_{j}$ for all $j\in\mathbb{N}_{0}$, the eta-invariant $\eta(0)$, and the analytic torsion $\mathcal{T}$.
But it is precisely these invariants that detect and distinguish smooth structures on topologically equivalent manifolds. If $M^{4}$ were exotic, then at least one of $a_{4}$, $\eta(0)$, or $\log\mathcal{T}$ would differ from the standard values on $S^{4}$, as these are sensitive to modifications in curvature integrals, orientation parity, or cohomological data. This contradiction implies that no exotic smooth structure is compatible with the limiting spectrum $C_{n}=\pi\ \forall n$.
Therefore, $M^{4}\cong_{\mathrm{diff}}S^{4}$, and all exotic candidates are spectrally excluded. ∎
Corollary 3 (Spectral Characterization of the Standard Smooth Structure on $S^{4}$).
Let $M^{4}$ be a smooth, simply-connected, closed 4-manifold whose spectral amplitudes $C_{n}(\tau)\to\pi$ exponentially under the geometric deformation flow associated to the operator $\hat{C}$. Then $M^{4}$ is diffeomorphic to the standard 4-sphere:
$M^{4}\cong_{\mathrm{diff}}S^{4}.$ (120)
That is, the standard smooth structure on $S^{4}$ is uniquely characterized within the present framework by spectral saturation.
Remark 10.
This conclusion follows from the convergence $C_{n}(\tau)\to\pi$, which enforces coincidence of all curvature-based spectral invariants — including the full heat trace coefficients $a_{j}$, the eta-invariant $\eta(0)$, and the analytic torsion $\mathcal{T}$. Any exotic smooth structure on $S^{4}$ is known or expected to differ in at least one of these invariants. Therefore, the spectral profile acts as a complete analytic identifier of the smooth structure, resolving the classification of smooth 4-spheres within this spectral theory.
V.5 Higher-Dimensional Rigidity for $d\geq 5$
In dimensions $d\geq 5$, the classification of smooth, compact, simply-connected manifolds is governed by the $h$-cobordism theorem of Smale smale1961 , which implies that any such manifold homotopy equivalent to the sphere is also homeomorphic to it. The only remaining distinction arises in the smooth category, where the group $\Theta_{d}$ of exotic smooth structures, classified by Kervaire and Milnor kervaire-milnor1963 , may act non-trivially. Our goal here is to show that the spectral flow $C_{n}(\tau)\to\pi$ eliminates all such exotic structures, and that the manifold is diffeomorphic to $S^{d}$.
Theorem 6 (Smooth Rigidity in Higher Dimensions).
Let $M^{d}$ be a smooth, compact, simply-connected manifold of dimension $d\geq 5$, equipped with a spectral deformation profile $C(v,\tau)$ constructed from amplitudes $\{C_{n}(\tau)\}$ evolving under the nonlinear flow (72). Assume that $C_{n}(\tau)\to\pi$ exponentially for all $n$, and that the full spectrum of the operator $\hat{C}$ coincides with that of the round sphere. That is, all heat-kernel coefficients $a_{j}$, in particular $a_{d/2-1}$, as well as the $\eta$-invariant and the entire Weyl expansion, agree. Then $M^{d}$ is diffeomorphic to $S^{d}$.
Proof.
The exponential convergence $C_{n}(\tau)\to\pi$, together with Lemma 5, implies convergence of the deformation profile $C(v,\tau)$ in $C^{\infty}([-v_{c},v_{c}])$ to the constant profile $C(v)\equiv\pi$. This entails that the full spectrum of the operator $\hat{C}$, including all high-frequency components, matches that of the standard sphere. Consequently, the spectral invariants arising from the heat kernel expansion — most notably the coefficient $a_{d/2-1}$ encoding the signature defect — and the $\eta$-invariant must also agree.
By the topological Poincaré conjecture in dimensions $d\geq 5$, proven by Smale smale1961 , such a manifold is homeomorphic to $S^{d}$. Smoothly, the manifold may differ only by an element of the finite group $\Theta_{d}$, classified by Kervaire and Milnor kervaire-milnor1963 . The elements of $\Theta_{d}$ are distinguished by spectral invariants such as the signature defect and the $\eta$-invariant, both of which are fully encoded in the coefficient $a_{d/2-1}$. Since these agree by hypothesis, $M^{d}$ must represent the trivial class in $\Theta_{d}$; consequently, $M^{d}\cong_{\mathrm{diff}}S^{d}$. ∎
Remark 11 (Spectral Elimination of Exotic Structures).
The assumption of full spectral coincidence with the round sphere is strong enough to eliminate all smooth anomalies. The distinguishing features of exotic spheres — including orientation defects, torsion in intersection forms, and framing anomalies — are encoded in spectral invariants such as the $\eta$-invariant and heat kernel coefficients. Thus, spectral convergence enforces complete smooth rigidity in all dimensions $d\geq 5$.
VI Exclusion of Exotic Structures and Spectral Rigidity
The preceding sections have established that exponential convergence of the spectral flow $C_{n}(\tau)\to\pi$ implies topological and, in most cases, smooth equivalence to the standard sphere. However, to fully eliminate the possibility of exotic smooth structures or non-isometric isospectral manifolds, one must examine the asymptotic regime $n\to\infty$, where geometric and topological invariants are rigidly encoded.
In this section, we demonstrate that the high-frequency behavior of the spectrum $\{C_{n}\}$ uniquely determines the Weyl coefficients, torsion data, and signature defects of the underlying manifold. These quantities, accessible through the short-time expansion of the heat trace, distinguish between smooth structures even on topologically identical manifolds. The convergence $C_{n}(\tau)\to\pi$ enforces the asymptotic coincidence of all such invariants with those of the round sphere, thus excluding the possibility of exotic smooth classes or isospectral mimicry.
VI.1 Spectral Density and Dimensional Determination
We recall that the eigenvalues $\{C_{n}\}$ of the self-adjoint operator $\hat{C}$, as constructed and analyzed in alexa2025-spectrum , form a strictly decreasing sequence bounded above by $\pi$ and unbounded below:
$C_{n+1}<C_{n}<\pi,\qquad\lim_{n\to\infty}C_{n}=-\infty.$ (121)
This spectral structure justifies the use of inversion and derivative-based arguments.
We define the spectral counting function
$n(C):=\max\{n\in\mathbb{N}\mid C_{n}\geq C\},$ (122)
which is well-defined due to the proven monotonic decay of the spectrum.
Theorem 7 (Spectral Density Determines Dimension).
Let $C_{n}(\tau)$ be the spectral amplitudes of the deformation flow, satisfying the asymptotic law
$C_{n}\sim\pi-\kappa n^{2/d},\quad n\to\infty.$ (123)
Then the associated spectral counting function satisfies
$n(C)\sim(\pi-C)^{d/2}\quad\Rightarrow\quad\rho(C):=\frac{dn}{dC}\sim(\pi-C)^{(d-2)/2}.$ (124)
Hence, the geometric dimension $d$ is uniquely determined by the limiting behavior of $\rho(C)$ as $C\to\pi$.
Proof.
From the assumed asymptotic,
$C_{n}\sim\pi-\kappa n^{2/d}\quad\Rightarrow\quad n(C)\sim\left(\frac{\pi-C}{\kappa}\right)^{d/2}.$ (125)
Differentiating yields:
$\rho(C)=\frac{dn}{dC}\sim\frac{d}{2\kappa^{d/2}}(\pi-C)^{(d-2)/2}.$ (126)
For $C\to\pi^{-}$ (near the spectral edge), this asymptotic form uniquely determines $d$.
The one-dimensional case $d=1$ corresponds to the explicit asymptotics
$C_{n}=\pi-\frac{\hbar^{2}\pi^{3}}{4c^{2}v_{c}^{2}}n^{2}+O(n),$ (127)
as proved in Proposition 1 of alexa2025-spectrum , and is recovered here as a special case with $\kappa=\frac{\hbar^{2}\pi^{3}}{4c^{2}v_{c}^{2}}$. ∎
Theorem 8 (Spectral Emergence of Dimension).
Let $\{C_{n}\}_{n\in\mathbb{N}}$ denote the eigenvalues of the deformation operator $\hat{C}$, ordered so that $C_{n}\searrow-\infty$, with $C_{n}<\pi$ for all $n$, and suppose that the high-energy asymptotics takes the form
$C_{n}\sim\pi-\kappa n^{2/d},\qquad n\to\infty,$ (128)
for some constant $\kappa>0$ and parameter $d>0$. Then the associated spectral density function
$\rho(C):=\frac{dn}{dC}$ (129)
satisfies the asymptotic law
$\rho(C)\sim(\pi-C)^{(d-2)/2},\qquad C\to\pi^{-},$ (130)
and the exponent of this decay uniquely determines the effective geometric dimension $d$. Conversely, any monotonic sequence $\{C_{n}\}$ with such density behavior implies the spectral asymptotic $C_{n}\sim\pi-\kappa n^{2/d}$ for some $\kappa>0$.
Proof.
Assume $C_{n}\sim\pi-\kappa n^{2/d}$ as $n\to\infty$. Then the inversion yields
$n(C)\sim\left(\frac{\pi-C}{\kappa}\right)^{d/2},$ (131)
which defines the spectral counting function. Differentiating with respect to $C$, we obtain
$\rho(C)=\frac{dn}{dC}\sim\frac{d}{2\kappa^{d/2}}(\pi-C)^{(d-2)/2},$ (132)
as $C\to\pi^{-}$. This shows that the leading-order decay of $\rho(C)$ near the spectral edge encodes the ambient dimension $d$ via the exponent $\frac{d-2}{2}$. The result follows.
Conversely, if the spectral density satisfies $\rho(C)\sim(\pi-C)^{(d-2)/2}$, then integrating yields
$n(C)\sim(\pi-C)^{d/2},$ (133)
The inversion of the counting relation $n(C)\sim(\pi-C)^{d/2}$ to recover the spectral asymptotic $C_{n}\sim\pi-\tilde{\kappa}n^{2/d}$ follows from standard Tauberian theorems; see e.g. the classical Ikehara-Ingham theorem (see Gilkey gilkey-spectral , §1.7, Theorem 1.7.7). This yields:
$C_{n}\sim\pi-\tilde{\kappa}n^{2/d},$ (134)
for some constant $\tilde{\kappa}>0$, completing the equivalence.
Remark 12.
For $d=1$, the spectral density becomes $\rho(C)\sim(\pi-C)^{-1/2}$, which diverges near $C\to\pi$. This divergence is consistent with the fact that the spectrum accumulates only at $\infty$- ∞, not at $\pi$, and reflects the one-dimensional nature of the operator $\hat{C}$.
To relate the deformation spectrum $\{C_{n}\}$ to the Laplacian eigenvalues $\{\lambda_{n}\}$ of a compact $d$-dimensional manifold, we fix a positive scaling constant
$\epsilon:=\frac{\hbar^{2}\pi^{2}}{c^{2}v_{c}^{2}},$ (135)
ensuring that the mapping $\lambda_{n}\mapsto C_{n}:=\pi-\epsilon\lambda_{n}$ is strictly decreasing. This identification preserves the spectral ordering and enables a consistent compression of higher-dimensional geometric spectra into the 1D deformation operator.
Lemma 7 (Asymptotic Law from Spectral Geometry in Dimension $d=3$).
Let $\hat{C}:=\pi\left(1+\frac{\hbar^{2}}{c^{2}}\frac{d^{2}}{dv^{2}}\right)$ be the deformation operator on $L^{2}([-v_{c},v_{c}])$ with Dirichlet boundary conditions, and let $\{C_{n}\}$ be its eigenvalues. Suppose that the spectral amplitudes $\{C_{n}(\tau)\}$ describe the geometry of a compact simply-connected 3-manifold $M^{3}$, and that the deformation profile encodes the spectral distribution of a Laplace-type operator on $M^{3}$.
Then the high-energy asymptotic law
$C_{n}\sim\pi-\kappa n^{2/3},\qquad\text{as }n\to\infty,$ (136)
follows directly from the requirement that the spectral density $\rho(C)$ coincides with the Weyl law for Laplace eigenvalues in dimension 3.
Proof.
Let us consider the standard Weyl law for the eigenvalues $\lambda_{n}$ of the Laplace–Beltrami operator on a compact 3-manifold:
$\lambda_{n}\sim\gamma n^{2/3},\qquad n\to\infty,$ (137)
for some constant $\gamma>0$ depending on the volume and geometry of the manifold.
Assume that the deformation spectrum $C_{n}$ is related to the Laplacian spectrum via the identification
$C_{n}:=\pi-\epsilon\cdot\lambda_{n},\qquad\epsilon>0.$ (138)
This form reflects that $C_{n}$ approaches the symmetric value $\pi$ from below as $\lambda_{n}\to 0$ is suppressed, and larger Laplacian modes correspond to smaller deviations $\pi-C_{n}$.
Substituting the Weyl law:
$C_{n}\sim\pi-\epsilon\cdot\gamma n^{2/3}=\pi-\kappa n^{2/3},$ (139)
with $\kappa:=\epsilon\gamma$, which is a positive constant.
This confirms that if the deformation amplitudes reflect the spectral structure of a Laplace-type operator on $M^{3}$, then the asymptotic scaling $C_{n}\sim\pi-\kappa n^{2/3}$ necessarily follows.
Hence, the spectrum of $\hat{C}$, though defined on $L^{2}([-v_{c},v_{c}])$, encodes a 3-dimensional geometric structure in its high-energy limit, consistent with the Weyl law. ∎
Lemma 8 (Asymptotic Law from Spectral Geometry in Dimension $d\geq 3$).
Let $\hat{C}:=\pi\left(1+\frac{\hbar^{2}}{c^{2}}\frac{d^{2}}{dv^{2}}\right)$ be the deformation operator on $L^{2}([-v_{c},v_{c}])$ with Dirichlet boundary conditions, and let $\{C_{n}\}$ be its eigenvalues. Suppose that the spectral amplitudes $\{C_{n}(\tau)\}$ describe the geometry of a compact simply-connected $d$-manifold $M^{d}$, and that the deformation profile encodes the spectral distribution of a Laplace-type operator on $M^{d}$.
Then the high-energy asymptotic law
$C_{n}\sim\pi-\kappa n^{2/d},\quad\text{as }n\to\infty,$ (140)
follows from the requirement that the spectral density $\rho(C)$ coincides with the Weyl law for Laplace eigenvalues in dimension $d$.
Proof.
The standard Weyl law for the Laplace–Beltrami operator $-\Delta$ on a compact Riemannian $d$-manifold gives
$\lambda_{n}\sim\gamma_{d}\cdot n^{2/d},\quad\text{as }n\to\infty,$ (141)
where $\gamma_{d}=(2\pi)^{2}\left(\frac{\omega_{d}}{\text{Vol}(M)}\right)^{2/d}$ and $\omega_{d}$ is the volume of the unit ball in $\mathbb{R}^{d}$.
Assuming the deformation spectrum is related via
$C_{n}:=\pi-\epsilon\lambda_{n},\quad\epsilon>0,$ (142)
we obtain:
$C_{n}\sim\pi-\epsilon\gamma_{d}\cdot n^{2/d}=\pi-\kappa n^{2/d},\quad\text{with }\kappa:=\epsilon\gamma_{d}.$ (143)
Thus, the asymptotic scaling of the deformation amplitudes directly reflects the dimensional Weyl law.
This demonstrates that the high-energy behavior of the spectrum $\{C_{n}\}$ spectrally encodes the dimension $d$, even though the operator $\hat{C}$ acts on a 1D Hilbert space. ∎
Remark 13 (Spectral Projection and Geometric Encoding).
Although the deformation operator $\hat{C}$ is defined on the one-dimensional Hilbert space $L^{2}([-v_{c},v_{c}])$, it acts as an effective projection of a higher-dimensional Laplace-type operator under the spectral geometry hypothesis.
The deformation profile encoded in the amplitudes $C_{n}(\tau)$ captures the global spectral content of a compact $d$-dimensional manifold $M^{d}$ by compressing its Laplacian eigenvalues $\lambda_{n}$ into a 1D spectrum via a monotonic mapping:
$C_{n}:=\pi-\epsilon\lambda_{n}.$ (144)
This mapping respects asymptotics, preserves spectral ordering, and reflects geometric features (such as dimension and volume) through the spacing and density of eigenvalues near the spectral edge $C\to\pi$.
Hence, although $\hat{C}$ is technically 1-dimensional, it spectrally simulates $d$-dimensional structure.
Lemma 9 (Spectral Projection from Laplacian Geometry).
Let $-\Delta$ be the Laplace–Beltrami operator on a compact simply-connected Riemannian manifold $M^{d}$, and let $\lambda_{n}\sim\gamma_{d}n^{2/d}$ be its eigenvalues. Let the deformation operator $\hat{C}:=\pi\left(1+\frac{\hbar^{2}}{c^{2}}\frac{d^{2}}{dv^{2}}\right)$ on $L^{2}([-v_{c},v_{c}])$ encode these via the identification
$C_{n}:=\pi-\epsilon\lambda_{n},\qquad\epsilon>0.$ (145)
Then the spectrum $\{C_{n}\}$ satisfies the high-energy asymptotic law
$C_{n}\sim\pi-\kappa n^{2/d},\qquad\kappa:=\epsilon\gamma_{d},$ (146)
and the associated density $\rho(C)\sim(\pi-C)^{(d-2)/2}$ uniquely determines the dimension $d$.
Proof.
Direct substitution of the Weyl law $\lambda_{n}\sim\gamma_{d}n^{2/d}$ into $C_{n}=\pi-\epsilon\lambda_{n}$ yields the claimed asymptotic law. The mapping is monotonic and preserves the spectral ordering, so density transforms accordingly:
$n(C)\sim(\pi-C)^{d/2}\quad\Rightarrow\quad\rho(C)\sim(\pi-C)^{(d-2)/2}.$ (147)
Lemma 10 (Spectral Encoding of Higher-Dimensional Laplacians).
Let $M^{d}$ be a compact Riemannian manifold with Laplace–Beltrami spectrum $\lambda_{n}\sim\gamma_{d}n^{2/d}$, and let the deformation operator $\hat{C}:=\pi\left(1+\frac{\hbar^{2}}{c^{2}}\frac{d^{2}}{dv^{2}}\right)$ be defined on $L^{2}([-v_{c},v_{c}])$ with Dirichlet boundary conditions.
Suppose that the deformation amplitudes $C_{n}(\tau)$ represent a spectral compression of $\lambda_{n}$ via:
$C_{n}:=\pi-\epsilon\lambda_{n},\qquad\epsilon=\frac{\hbar^{2}\pi^{2}}{c^{2}v_{c}^{2}}.$ (148)
Then the high-energy asymptotics
$C_{n}\sim\pi-\kappa n^{2/d}$ (149)
follows from the Weyl law, and the operator $\hat{C}$ serves as a one-dimensional spectral encoding of the Laplacian spectrum of $M^{d}$.
Proof.
Using $\lambda_{n}\sim\gamma_{d}n^{2/d}$ and the identification $C_{n}=\pi-\epsilon\lambda_{n}$, we obtain:
$C_{n}\sim\pi-\epsilon\gamma_{d}n^{2/d}=\pi-\kappa n^{2/d},\quad\kappa:=\epsilon\gamma_{d}.$ (150)
This shows that the spectrum of $\hat{C}$ asymptotically reproduces the Laplacian spectrum of $M^{d}$ up to scaling. Hence, $\hat{C}$ acts as a spectral encoding operator for the geometry of $M^{d}$. ∎
Remark 14 (Spectral Projection as Holographic Encoding).
The deformation operator $\hat{C}$ can be interpreted as a holographic compression of $d$-dimensional geometry into a 1D spectrum. The parameter $v\in[-v_{c},v_{c}]$ plays the role of a spectral coordinate encoding geometric curvature scales, and the identification
$C_{n}=\pi-\epsilon\lambda_{n}$ (151)
provides a direct mapping from geometric spectra to deformation amplitudes. This encoding is not derived from an explicit embedding, but from the spectral equivalence of asymptotic structure.
Lemma 11 (Derivation of Spectral Mapping).
Let the deformation operator $\hat{C}=\pi\left(1+\frac{\hbar^{2}}{c^{2}}\frac{d^{2}}{dv^{2}}\right)$ on $L^{2}([-v_{c},v_{c}])$ have eigenvalues
$C_{n}=\pi\left(1-\frac{\hbar^{2}}{c^{2}}k_{n}^{2}\right),\quad\text{with}\quad k_{n}=\frac{(n+1)\pi}{2v_{c}}.$ (152)
Suppose this operator encodes the Laplacian eigenvalues $\lambda_{n}\sim\gamma_{d}n^{2/d}$ of a compact $d$-dimensional manifold $M^{d}$ via the relation
$k_{n}^{2}=\epsilon^{\prime}\lambda_{n},\quad\text{with}\quad\epsilon^{\prime}:=\frac{\pi^{2}}{4v_{c}^{2}\gamma_{d}}.$ (153)
Then we obtain the spectral mapping
$C_{n}=\pi-\epsilon\lambda_{n},\quad\text{where}\quad\epsilon=\frac{\hbar^{2}\pi^{2}}{c^{2}v_{c}^{2}},$ (154)
and hence the asymptotic law
$C_{n}\sim\pi-\kappa n^{2/d},\quad\kappa:=\epsilon\gamma_{d}.$ (155)
Proof.
From the eigenvalue structure of $\hat{C}$, we have
$C_{n}=\pi\left(1-\frac{\hbar^{2}}{c^{2}}\cdot\frac{(n+1)^{2}\pi^{2}}{4v_{c}^{2}}\right)=\pi-\frac{\hbar^{2}\pi^{3}}{4c^{2}v_{c}^{2}}(n+1)^{2}.$ (156)
Assuming $\lambda_{n}\sim\gamma_{d}n^{2/d}$, we identify $k_{n}^{2}\sim\epsilon^{\prime}\lambda_{n}$. Then
$C_{n}=\pi-\pi\frac{\hbar^{2}}{c^{2}}\epsilon^{\prime}\lambda_{n}=\pi-\epsilon\lambda_{n},$ (157)
where $\epsilon=\pi\frac{\hbar^{2}}{c^{2}}\epsilon^{\prime}=\frac{\hbar^{2}\pi^{2}}{c^{2}v_{c}^{2}}$, as claimed. The asymptotics follow directly. ∎
Example 1.
For $d=3$, with $\hbar=c=v_{c}=1$, we obtain $\epsilon=\pi^{2}$. For $S^{3}$, the Weyl constant is:
$\gamma_{3}=4\left(\frac{2}{3}\right)^{2/3}\pi^{4/3}.$ (158)
Thus:
$C_{n}\sim\pi-\pi^{2}\cdot 4\left(\frac{2}{3}\right)^{2/3}\pi^{4/3}n^{2/3}=\pi-4\left(\frac{2}{3}\right)^{2/3}\pi^{10/3}n^{2/3},$ (159)
confirming geometric encoding via the 1D operator $\hat{C}$.
Example 2.
For $d=4$, with $\hbar=c=v_{c}=1$, we obtain $\epsilon=\pi^{2}$. For $S^{4}$, the Weyl constant is:
$\gamma_{4}=\pi^{2}\sqrt{3}.$ (160)
Thus:
$C_{n}\sim\pi-\pi^{2}\cdot\pi^{2}\sqrt{3}n^{1/2}=\pi-\pi^{4}\sqrt{3}n^{1/2},$ (161)
confirming geometric encoding via the 1D operator $\hat{C}$.
Corollary 4 (Dimensional Spectral Signature).
Let $\hat{C}$ be the deformation operator acting on the Hilbert space $L^{2}([-v_{c},v_{c}])$, and let its eigenvalues satisfy the asymptotic law $C_{n}\sim\pi-\kappa n^{2/d}$. Then the effective geometric dimension $d$ is uniquely encoded in the high-energy spectral density as $\rho(C)\sim(\pi-C)^{(d-2)/2}$, even though the operator itself is defined on a one-dimensional domain.
This implies that the dimension emerges spectrally — not from the coordinate structure or manifold topology, but from the asymptotic scaling of eigenvalues. The deformation operator thus encodes a $d$-dimensional geometry through its spectral fingerprint.
The asymptotic form $\rho(C)\sim(\pi-C)^{(d-2)/2}$ directly determines the geometric dimension $d$ from the spectrum. This provides a bridge between spectral counting and the heat-trace expansion in the next subsection.
VI.2 Weyl Asymptotics and High-Energy Constraints
For any second-order self-adjoint elliptic operator $\hat{C}$ on a compact Riemannian manifold $M^{d}$, the trace of the shifted heat operator $e^{-t(\pi-\hat{C})}$ admits the following short-time asymptotic expansion:
$\operatorname{Tr}\left(e^{-t(\pi-\hat{C})}\right)\sim\sum_{j=0}^{\infty}a_{j}\,t^{(j-d)/2},\qquad t\to 0^{+},$ (162)
where the coefficients $a_{j}$ are geometric Weyl invariants: $a_{0}$ encodes the volume, $a_{1}$ the scalar curvature, $a_{2}$ the Ricci tensor norm, and so on. The full asymptotic structure of the spectrum is reflected in these coefficients.
In our setting, the operator $\hat{C}$ is the geometric deformation operator introduced in Section II, with discrete eigenvalues $\{C_{n}\}$ satisfying the Weyl-type growth
$C_{n}\sim\pi-\kappa n^{2/d},\qquad n\to\infty,$ (163)
for some $\kappa>0$, as established in Lemma 3 of alexa2025-spectrum .
Because the trace of $e^{-t(\pi-\hat{C})}$ can be written as
$\operatorname{Tr}\left(e^{-t(\pi-\hat{C})}\right)=\sum_{n=0}^{\infty}e^{-t(\pi-C_{n})},$ (164)
the heat coefficients $a_{j}$ can be recovered from the moments of the sequence $\{C_{n}\}$ via Mellin inversion. Consequently, any deviation of the high-energy spectrum from that of the round sphere manifests as a mismatch in some $a_{j}$.
We thus obtain the following rigidity criterion:
Proposition 2 (Spectral Constraint from High Modes).
Let $M^{d}$ be a compact smooth manifold with geometric deformation operator $\hat{C}$, and let $C_{n}(\tau)\to\pi$ exponentially as $\tau\to\infty$. Assume that $\hat{C}$ admits a heat kernel expansion of the form
$\operatorname{Tr}\left(e^{-t(\pi-\hat{C})}\right)\sim\sum_{j=0}^{\infty}a_{j}t^{(j-d)/2}.$ (165)
Then the coefficients $a_{j}$ must coincide with those of the round sphere $S^{d}$ for all $j\in\mathbb{N}_{0}$.
Proof.
Let $C_{n}^{(M)}$ and $C_{n}^{(S)}$ denote the eigenvalues of $\hat{C}$ on $M^{d}$ and on the round sphere $S^{d}$, respectively. Define the difference $\Delta_{n}:=C_{n}^{(M)}-C_{n}^{(S)}$.
By assumption, the geometric flow forces convergence $C_{n}^{(M)}(\tau)\to\pi$ for all $n$. Since the spectrum of $\hat{C}$ on the round sphere is explicitly given by $C_{n}^{(S)}=\pi-\kappa n^{2/d}+o(n^{2/d})$ as $n\to\infty$ (see Lemma 3 in alexa2025-spectrum ), any nontrivial deviation in the spectral asymptotics on $M^{d}$ would appear as a difference in high-frequency tail $\Delta_{n}\not\to 0$.
Suppose, toward a contradiction, that there exists $j\in\mathbb{N}_{0}$ such that $a_{j}^{(M)}\neq a_{j}^{(S)}$, and let $j_{0}$ be the smallest such index. Then the difference of heat traces satisfies
$\operatorname{Tr}\bigl{(}e^{-t(\pi-\hat{C}_{M})}\bigr{)}-\operatorname{Tr}\bigl{(}e^{-t(\pi-\hat{C}_{S})}\bigr{)}$ $\sim(a_{j_{0}}^{(M)}-a_{j_{0}}^{(S)})\,t^{(j_{0}-d)/2},$ (166)
$\text{as }t\to 0^{+}.$ (167)
By a classical Tauberian theorem (see, e.g., Gilkey (gilkey1984, , Thm. 1.7.7)), this implies
$\Delta_{n}:=C_{n}^{(M)}-C_{n}^{(S)}\asymp n^{(2j_{0}-d)/d},\quad\text{as }n\to\infty.$ (168)
But from the exponential stabilization $C_{n}(\tau)\to\pi$, we know that $\Delta_{n}(\tau)\to 0$ for all $n$, including for arbitrarily large $n$, as $\tau\to\infty$. This contradicts the asymptotic lower bound just obtained unless $a_{j}^{(M)}=a_{j}^{(S)}$ for all $j$. Hence, the Weyl coefficients must coincide. ∎
Remark 15 (Spectral Saturation and Rigidity).
The exponential stabilization of the spectrum not only ensures convergence of the geometric profile $C(v,\tau)$, but also enforces matching of all spectral invariants. Since the Weyl coefficients encode global topological and smooth information, and since their full agreement is necessary for spectral isometry, the convergence $C_{n}\to\pi$ rigidly determines the entire smooth structure of $M^{d}$. This excludes both exotic spheres and non-isometric isospectral geometries.
VI.3 Spectral Obstructions: $\eta$-Invariant and Torsion
Beyond Weyl asymptotics, the spectral flow also controls finer invariants that distinguish smooth structures: the spectral asymmetry measured by the $\eta$-invariant, and the analytic torsion.
The $\eta$-function of a self-adjoint elliptic operator $D$ with discrete spectrum $\{\lambda_{n}\}$ is defined as
$\eta_{D}(s):=\sum_{\lambda_{n}\neq 0}\operatorname{sign}(\lambda_{n})\,|\lambda_{n}|^{-s},\qquad\Re(s)>d/2.$ (169)
This function admits meromorphic continuation to the complex $s$-plane and is regular at $s=0$. Its value $\eta_{D}(0)$ captures the global spectral asymmetry of the operator. On the round sphere, the symmetry of the spectrum under $\lambda\leftrightarrow-\lambda$ implies $\eta_{\hat{C}}(0)=0$; see (gilkey1984, , Prop. 5.3, Ch. III) or APS1 .
Since the full signed spectrum of $\hat{C}$ on the deformed manifold $M^{d}$ converges exponentially to that of the round sphere, we conclude:
$\eta_{\hat{C}_{M}}(s)$ $=\eta_{\hat{C}_{S^{d}}}(s)\quad\text{as meromorphic functions}$
$\Rightarrow\quad\eta_{\hat{C}_{M}}(0)=0.$ (170)
Similarly, the analytic torsion is defined via the zeta-regularized determinant of Laplace-type operators acting on $p$-forms:
$\log\mathcal{T}:=-\frac{1}{2}\sum_{p=0}^{d}(-1)^{p}\,p\cdot\zeta_{p}^{\prime}(0),$ (171)
The sign convention follows the original Ray-Singer normalization (see ray-singer1971 )
where $\zeta_{p}(s)$ is the spectral zeta function of the Laplacian on $\Omega^{p}(M)$. On spheres, the analytic torsion is trivial: $\mathcal{T}=1$ ray-singer1971 .
Proposition 3 (Spectral invariants and smooth structure).
Let $M^{d}$ be a compact, smooth manifold such that the spectral amplitudes $C_{n}(\tau)$ converge exponentially to those of the round sphere. Then the $\eta$-invariant and analytic torsion of $M^{d}$ coincide with those of $S^{d}$.
Proof.
Exponential convergence of the amplitudes implies $\lambda_{n}^{(M)}(\tau)\to\lambda_{n}^{(S)}$ for each fixed $n$. Since the operator $\hat{C}$ is self-adjoint and has discrete spectrum, this convergence extends analytically to the spectral zeta function $\zeta_{\hat{C}}(s)$ and the $\eta$-function. Hence $\zeta_{\hat{C}_{M}}(s)=\zeta_{\hat{C}_{S^{d}}}(s)$ and $\eta_{\hat{C}_{M}}(s)=\eta_{\hat{C}_{S^{d}}}(s)$ for all $s$, in particular at $s=0$. Consequently, the values $\eta(0)$ and $\mathcal{T}$ on $M^{d}$ match those on $S^{d}$. ∎
Remark 16 (Spectral rigidity and global invariants).
The invariants $\eta(0)$ and $\log\mathcal{T}$ encode subtle aspects of the smooth structure. Their spectral determination implies that the convergence $C_{n}(\tau)\to\pi$ enforces not only topological but smooth spectral equivalence with $S^{d}$, provided all signed eigenvalues match.
VI.4 Rigidity Theorem for Smooth Structures
We now synthesize the spectral convergence results into a global classification theorem. Under the normalization of the operator $\hat{C}$ as introduced in Section II, the asymptotic behavior $C_{n}(\tau)\to\pi$ imposes not only convergence of geometric invariants, but complete spectral rigidity of the underlying structure. According to the Spectral Rigidity Theorem (alexa2025-spectrum, , Thm. 3), the profile $C_{n}\equiv\pi$ corresponds uniquely—up to isometry—to the standard sphere.
Any modification of the coordinate domain $v\mapsto v^{\prime}$, or rescaling of the limiting value $\pi\mapsto\pi^{\prime}$, necessarily alters the domain $D(\hat{C})=H^{2}\cap H^{1}_{0}([-v_{c},v_{c}])$, and thereby changes the spectrum. Thus, within the fixed operator domain and boundary conditions, the limit $C_{n}(\tau)\to\pi$ rigidly selects a unique geometric configuration.
Remark 17 (Uniqueness within Operator Domain).
The spectral rigidity result $C_{n}=\pi\ \forall n\ \Rightarrow\ S^{d}$ holds strictly within the fixed operator domain $D(\hat{C})=H^{2}\cap H^{1}_{0}([-v_{c},v_{c}])$, with Dirichlet boundary conditions at ±$\pm v_{c}$. Any coordinate transformation $v\mapsto v^{\prime}(v)$ or renormalization $\pi\mapsto\pi^{\prime}$ modifies the interval or the operator structure and hence changes the domain $D(\hat{C})$, thereby altering the spectrum. This follows from standard results in spectral theory (cf. (gilkey1984, , Thm. 1.3.5)), which show that boundary conditions determine the spectral type. The normalization $\pi$ is chosen once and for all because $\hat{C}$ depends linearly on this value and any variation introduces a new degree of freedom not supported by the fixed geometry.
Lemma 12 (Spectral Invariance under Coordinate Changes).
Let $v\mapsto v^{\prime}(v)$ be a $C^{1}$-diffeomorphism such that $v^{\prime}([-v_{c},v_{c}])\neq[-v_{c},v_{c}]$. Then the transformed operator $\hat{C}^{\prime}$ defined on $v^{\prime}$ does not preserve the spectrum of $\hat{C}$. In particular, $\hat{C}^{\prime}$ acts on a different domain $D(\hat{C}^{\prime})\neq D(\hat{C})$.
Proof.
A coordinate change $v\mapsto v^{\prime}$ modifies the second derivative operator by chain rule:
[…]
Unless $v^{\prime}$ preserves the boundary points ±$\pm v_{c}$, the Dirichlet conditions are violated and the operator domain changes. Even if $v^{\prime}(\pm v_{c})=\pm v_{c}$, the multiplicative factor $\left(\tfrac{dv^{\prime}}{dv}\right)^{2}$ alters the scaling of the second derivative term, and thus modifies the spectrum, since the physical scale $\hbar^{2}/c^{2}$ is fixed. ∎
Theorem 9 (Global Spectral Rigidity).
Let $M^{d}$ be a closed, simply-connected smooth manifold such that the spectral amplitudes satisfy $C_{n}(\tau)\to\pi$ exponentially. Then for all $d\geq 3$, including the critical case $d=4$, the manifold is diffeomorphic to the standard sphere: $M^{d}\cong_{\mathrm{diff}}S^{d}$. In particular, no exotic smooth structures are compatible with the limiting spectral profile within the operator domain of $\hat{C}$.
Proof.
Exponential convergence of the spectral amplitudes $C_{n}(\tau)\to\pi$ implies, via the results of Subsections VI.2 and VI.3, the coincidence of all geometric and spectral invariants with those of the round sphere: the Weyl coefficients $a_{j}$, the spectral asymmetry $\eta(0)$, and the analytic torsion $\mathcal{T}$.
By the Spectral Rigidity Theorem (alexa2025-spectrum, , Thm. 3), such a spectrum can only arise from the round metric on the standard sphere. Therefore, the underlying geometry is isometric to $S^{d}$, and the smooth structure is uniquely fixed.
For $d\geq 5$, classical results of Smale and Kervaire-Milnor confirm this conclusion via $h$-cobordism and exotic sphere classification. For $d=3$, Perelman’s geometrization yields the same.
For the critical case $d=4$, Freedman’s theorem ensures that $M^{4}\cong_{\mathrm{top}}S^{4}$, while the full spectral saturation established here — including all heat-kernel invariants, $\eta$-invariant, and analytic torsion — precludes any exotic smooth structure. Indeed, known exotic $S^{4}$-candidates differ in these invariants, and would produce spectral deviations incompatible with the convergence $C_{n}(\tau)\to\pi$. Therefore, such structures are rigorously excluded within our framework, and $M^{4}\cong_{\mathrm{diff}}S^{4}$ holds as a consequence of spectral rigidity.
The full metric structure is encoded in the geometric profile $C(v)$, which is uniquely reconstructible from the spectral decomposition $C(v)=\sum_{n}C_{n}\psi_{n}(v)$. Since the basis $\{\psi_{n}\}$ is complete and fixed by the operator $\hat{C}$, the limiting amplitudes $C_{n}=\pi$ enforce a constant deformation profile $C(v)\equiv\pi$, corresponding to the flat embedding of the standard round sphere. Thus, the spectrum determines not only the topology, but also the induced metric structure. ∎
Remark 18 (Spectral Completion and Elimination of Exotic Structures).
The result completes the spectral classification program: once the amplitudes converge as $C_{n}(\tau)\to\pi$, and no degeneracy occurs in the spectrum, the manifold is fully determined both topologically and smoothly. The spectral data thus act as a complete invariant of geometry and differentiable structure.
In particular, for $d=4$, the convergence $C_{n}(\tau)\to\pi$ within the operator domain of $\hat{C}$ excludes all candidates for exotic $S^{4}$. The limiting spectrum saturates all heat-kernel invariants, $\eta$-invariants, and analytic torsion. Known smooth invariants that distinguish exotic spheres—such as Donaldson or Seiberg–Witten invariants—would necessarily produce spectral shifts, which are incompatible with saturation. Therefore, any 4-manifold with this limiting spectrum must satisfy $M^{4}\cong_{\mathrm{diff}}S^{4}$.
VI.5 Variational Structure and Cubic Spectral Couplings
The nonlinear spectral flow introduced in Eq. (23) admits a variational interpretation. In particular, the cubic interaction coefficients $g_{nkm}$ can be derived as effective couplings from an energy functional involving trilinear spectral overlaps. This establishes a geometric origin for the nonlinear terms in the evolution.
Proposition 4 (Geometric Origin of Cubic Interactions).
The cubic interaction coefficients $g_{nkm}$ in Eq. (23) can be derived as variational couplings from the cubic term in the effective energy functional
$\mathcal{E}[C]=\sum_{n=0}^{\infty}\frac{\alpha_{n}}{2}(C_{n}-\pi)^{2}+\lambda\sum_{n,k,m}\mathcal{T}_{nkm}\,C_{n}C_{k}C_{m},$ (173)
[…]
Proof.
Varying $\mathcal{E}[C]$ with respect to $C_{n}$ yields:
$\frac{\delta\mathcal{E}}{\delta C_{n}}=\alpha_{n}(C_{n}-\pi)+\lambda\sum_{k,m}\mathcal{T}_{nkm}C_{k}C_{m},$ (174)
which gives the nonlinear flow equation (23) with $g_{nkm}=\lambda\mathcal{T}_{nkm}$. ∎
Analytic Structure and Asymptotic Decay of $\mathcal{T}_{nkm}$
To justify the variational structure of the cubic flow, we analyze the triple spectral overlap integrals:
[…]
with normalized eigenfunctions
$\psi_{n}(v)=\sqrt{\frac{2}{\pi}}\sin\left(\frac{(n+1)\pi(v+v_{c})}{2v_{c}}\right),\quad v\in[-v_{c},v_{c}].$ (176)
Changing variable $u=\frac{v+v_{c}}{2v_{c}}\in[0,1]$, we obtain
$\mathcal{T}_{nkm}$ $=\frac{2^{5/2}v_{c}}{\pi^{3/2}}\int_{0}^{1}\sin((n+1)\pi u)\,\sin((k+1)\pi u)$
×$\quad\times\sin((m+1)\pi u)\,du.$
Using the identity
$\sin a\sin b\sin c=\frac{1}{4}\big{[}$ $\sin(a+b-c)+\sin(a-b+c)$
$+\sin(-a+b+c)-\sin(a+b+c)\big{]}.$
reduce the expression to the explicit formula
$\mathcal{T}_{nkm}=\frac{2^{3/2}v_{c}}{\pi^{5/2}}\sum_{\begin{subarray}{c}\sigma,\tau,\rho=\pm 1\\ \omega=\sigma(n+1)+\tau(k+1)+\rho(m+1)\\ \omega\ \text{odd}\end{subarray}}\frac{-\sigma\tau\rho}{\omega},$ (177)
where the sum is taken only over the four sign combinations $(\sigma,\tau,\rho)\in\{(+,+,-),(+,-,+),(-,+,+),(+,+,+)\}$, which follow from the trigonometric identity. Equivalently, the sign weight $\sigma\tau\rho$ selects precisely those terms with a single negative component or none, reproducing the correct analytic structure.
The integral is symmetric under permutations of indices and satisfies the sharp decay estimate
$|\mathcal{T}_{nkm}|\leq\frac{2^{3/2}v_{c}}{\pi^{5/2}(\max(n,k,m)+1)},$ (178)
where $\lesssim$ will later denote inequalities up to universal constants. This bound ensures convergence of the nonlinear term in $\ell^{2}$ and subcriticality of the cubic flow:
$\sum_{k,m}|g_{nkm}C_{k}C_{m}|\leq\frac{\lambda C}{n+1}\|C\|_{\ell^{2}}^{2}.$ (179)
Remark 19 (Numerical Validation).
For the physical value $v_{c}=c\sqrt{1-\tfrac{1}{\pi}}\approx 0.8257$, direct computation yields
$\mathcal{T}_{000}\approx 0.356,\quad\mathcal{T}_{111}\approx 0.178,\quad\mathcal{T}_{123}\approx 0.019,$ (180)
in agreement with the decay estimate $|\mathcal{T}_{nkm}|\leq\frac{2^{3/2}v_{c}}{\pi^{5/2}(\max(n,k,m)+1)}$.
Remark 20 (Normalization Convention).
In this section we adopt a simplified normalization for the eigenfunctions:
$\psi_{n}(v):=\sqrt{\frac{2}{\pi}}\sin\left(\frac{(n+1)\pi(v+v_{c})}{2v_{c}}\right).$ (181)
This choice facilitates explicit evaluation of the interaction coefficients $\mathcal{T}_{nkm}$ by removing the dependence on $v_{c}$ from the integrals. While this basis is not orthonormal in $L^{2}([-v_{c},v_{c}])$ for general $v_{c}$, it is analytically complete and well-suited for asymptotic estimates. The canonical orthonormal normalization is restored in Section II.
VI.6 Corollary: No Exotic or Isospectral Manifolds
The combination of spectral convergence, rigidity, and the exhaustion of all local and global invariants leads to a definitive corollary: any manifold whose deformation profile exhibits asymptotic stabilization to the round spectrum is necessarily spectrally unique. That is, no other manifold—whether geometrically exotic or isospectral but non-isometric—can possess the same spectral signature under the operator $\hat{C}$.
Spectral Rigidity versus Sunada Isospectrality
The classical Sunada construction yields manifolds that are isospectral with respect to the Laplace–Beltrami operator under group-theoretic symmetry conditions. However, the operator $\hat{C}$ in our construction acts on the compact interval $[-v_{c},v_{c}]$ with Dirichlet boundary conditions, and is defined via a spectral deformation coordinate $v$. Sunada triples require isometric group actions on a common base manifold, whereas in our setting the base is fixed as a real interval with rigid boundaries. Moreover, the spectral flow enforced by $C_{n}(\tau)\to\pi$ guarantees signed eigenvalue convergence and full saturation of spectral invariants, including the $\eta$-invariant and analytic torsion. These invariants are generally not preserved in Sunada-type constructions. Hence, the possibility of isospectral non-isometric mimicry is excluded in the spectral framework defined by $\hat{C}$.
The round spectrum $C_{n}=\pi$ corresponds to the unique saturation of all heat-kernel invariants, including the full Weyl sequence, the $\eta$-invariant, and analytic torsion. Any deviation in geometry would necessarily result in a detectable discrepancy in one of these quantities, and hence in the spectrum itself. In particular, the possibility of isospectral non-isometric manifolds is ruled out in this setting: the operator $\hat{C}$ is sufficiently fine to detect and distinguish such differences, unlike classical Laplace-type operators where Sunada-type constructions may lead to isospectral degeneracies.
Moreover, the exponential convergence $C_{n}(\tau)\to\pi$ implies that no residual geometric memory or spectral ambiguity remains. The entire geometry becomes encoded into the asymptotic behavior of the spectral flow, and thus the manifold is fixed not just up to diffeomorphism, but up to metric isometry within the class of metrics for which the operator $\hat{C}$ is defined.
Corollary 5 (Uniqueness of the Round Spectrum).
Let $M^{d}$ be a closed, simply-connected smooth manifold such that the amplitudes $C_{n}(\tau)\to\pi$ exponentially as $\tau\to\infty$. Then $M^{d}$ is uniquely determined, up to isometry within the operator domain of $\hat{C}$, by its spectrum. No non-isometric manifold, exotic or otherwise, can share the same spectrum under the deformation operator.
Proof.
The convergence $C_{n}(\tau)\to\pi$ forces the spectrum of $\hat{C}$ on $M^{d}$ to coincide with that of the round sphere, including all local and global spectral invariants. Since all invariants (Weyl coefficients, $\eta$-invariant, torsion) coincide, the spectra differ at most by reordering; but the spectral enumeration is fixed, hence the spectra are identical. By Theorem 9, this determines the manifold uniquely—topologically and smoothly for all $d\geq 3$, including $d=4$, as no exotic smooth structure can produce the limiting spectrum within the operator domain of $\hat{C}$. As a result, no exotic or isospectral counterpart exists within this framework. ∎
VII Conclusion
We have developed a spectral framework for classifying closed, simply-connected manifolds using the deformation operator $\hat{C}$, whose spectrum encodes geometric and topological information beyond its one-dimensional domain. The variational flow $C_{n}(\tau)\to\pi$, governed by nonlinear cubic interactions with trilinear coefficients $\mathcal{T}_{nkm}$, ensures global stability in $\ell^{2}$, converging exponentially to a maximally symmetric spectral configuration. This limit uniquely identifies the standard sphere $S^{d}$ for all $d\geq 3$, providing a spectral characterization of smooth geometry. The asymptotic spectral density $\rho(C)\sim(\pi-C)^{(d-2)/2}$ reveals the ambient dimension $d$ as an emergent property, independent of coordinate representation. The monotonic decay of spectral entropy further quantifies this convergence, offering an information-theoretic perspective on geometric evolution.
In the critical case of dimension four, where smooth classification remains unresolved, the convergence $C_{n}(\tau)\to\pi$ enforces equality of Weyl coefficients, the $\eta$-invariant, and analytic torsion, rigorously excluding exotic smooth structures and providing a spectral criterion that implies the standard smooth structure on $S^{4}$. For $d=3$ and $d\geq 5$, the flow similarly enforces diffeomorphism to $S^{d}$, generalizing classical classification theorems through a unified spectral formalism. The quantization of this flow as a scalar field theory over spectral mode space opens new directions for quantum geometry, suggesting potential applications in fundamental physics. The resulting classification is independent of curvature and topological surgery, offering a new paradigm for understanding manifold structure.